### v1

Abstract
We present sharp estimates for the extremal eigenvalues of the Schur complements arising in saddle point problems. These estimates are derived using the auxiliary space theory, in which a given iterative method is interpreted as an equivalent but more elementary iterative method on an auxiliary space, enabling us to obtain sharp convergence estimates. The proposed framework improves or refines several existing results, which can be recovered as corollaries of our results. To demonstrate the versatility of the framework, we present various applications from scientific computing: the augmented Lagrangian method, mixed finite element methods, and nonoverlapping domain decomposition methods. In all these applications, the condition numbers of the corresponding Schur complements can be estimated in a straightforward manner using the proposed framework.
keywords:
Saddle point problems, Schur complements, Preconditioners, Auxiliary space theory, Nonoverlapping domain decomposition methods
{AMS}
65F10, 65N22, 65N55
1 Introduction
This paper is concerned with iterative methods for solving the general linear saddle point problem defined on finite-dimensional vector spaces $V$ and $W$:
(1) $\begin{bmatrix}A&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ p\end{bmatrix}=\begin{bmatrix}f\\ g\end{bmatrix}$,
where $A\colon V\to V$ is a linear operator that is either symmetric positive definite (SPD) or symmetric positive semidefinite (semi-SPD), $B\colon V\to W$ is a surjective linear operator, and $f\in V$, $g\in W$ are given data. We denote by $(u,p)$ the solution of (1). It is well-known that the system (1) is equivalent to the constrained quadratic optimization problem
(2) $\min_{v\in V}\left\{\frac{1}{2}(Av,v)-(f,v)\right\}\quad\text{subject to}\quad Bv=g,$
where $p$ serves as the Lagrange multiplier enforcing the constraint. Such saddle point problems arise in many areas of scientific computing; see [2] and the references therein.
There are three principal approaches for iterative methods for solving the saddle point problem (1[2, 41]. The first is to apply iterative methods to the SPD dual problem of (1) obtained in terms of the Schur complement. The second is to apply stationary iterations, such as Uzawa-type methods [4] or augmented Lagrangian methods [12]. The third is to apply Krylov subspace methods, such as MINRES [31], directly to (1) with suitable preconditioners [28]. In all three approaches, the properties of the Schur complement play a central role in both the design and the analysis of algorithms.
The purpose of this paper is to present sharp estimates for the extremal eigenvalues of the Schur complements arising in saddle point problems of the form (1). If AA is SPD, then the Schur complement $S=BA^{-1}B^{t}$ has the structure of an auxiliary space preconditioner [29, 40], where $A^{-1}$ corresponds to an iterative method defined on the auxiliary space $V$, and $B$ corresponds to the transfer operator mapping the auxiliary space $V$ to the original space $W$. A similar argument also applies in the case when AA is semi-SPD. Thus, the Schur complement $S$ can be analyzed by means of the recently proposed auxiliary space theory [33], which provides sharp estimates for iterative methods for both SPD and semi-SPD systems in a unified framework.
Estimates for the extremal eigenvalues of the Schur complement when AA is SPD are presented in Theorem 4.1, which are sharp in the sense that they are expressed as identities. Sharp estimates for the Schur complement preconditioned by a general SPD preconditioner are given in Theorem 4.5. These results refine existing abstract estimates for the Schur complement (see, e.g., [3, 25, 37]) in the sense that the earlier results can be recovered as straightforward corollaries. While our estimates follow directly from elementary linear algebra within the auxiliary space framework, the existing results required more involved arguments. Finally, our results extend naturally to the case when AA is semi-SPD, as shown in Theorems 4.9 and 4.10.
To demonstrate the usefulness of the proposed estimates for the Schur complement, we present several applications from scientific computing. The first example is the augmented Lagrangian method [12], a stationary iterative method for solving (1). Using the proposed framework, we provide sharp estimates for the error propagation operator of the augmented Lagrangian method. In particular, we recover the results of [17, 18, 21] showing that the convergence rate of the augmented Lagrangian method can be made arbitrarily fast by choosing the augmented Lagrangian parameter $\epsilon$ sufficiently small. The second example is mixed finite element methods for solving mixed variational problems [3]. We show how the proposed framework yields simplified analyses of Schur complements arising in mixed finite element discretizations of Darcy flow [5, 32] and the Stokes equations [1, 38]. The final example concerns nonoverlapping domain decomposition methods [37, 43], specifically FETI (finite element tearing and interconnecting) [11] and FETI-DP (dual–primal FETI) [9, 16]. For both unpreconditioned and preconditioned versions of FETI and FETI-DP, we provide simplified convergence analyses using the proposed framework.
The remainder of this paper is organized as follows. In Section 2, we briefly summarize the auxiliary space theory. In Section 3, we review well-posedness, Schur complement systems, and strategies for designing iterative methods for saddle point problems. In Section 4, we present sharp estimates for the extremal eigenvalues of the Schur complements. In Sections 5, 6, and 7, we demonstrate applications of our results to problems arising in scientific computing, including the augmented Lagrangian method, mixed finite element methods, and nonoverlapping domain decomposition methods. Finally, in Section 8, we conclude with some remarks.
1.1 Notation
We summarize here the notation used throughout the paper. Let $V$ and $W$ be finite-dimensional vector spaces equipped with inner products $(\cdot,\cdot)$ and the associated norms $\|\cdot\|$. For any subspace $S\subset V$, we denote by $S^{\perp}$ its orthogonal complement. For a linear operator $B\colon V\to W$, the null space and the range are denoted by $\mathcal{N}(B)$ and $\mathcal{R}(B)$, respectively, and the adjoint of $B$ is denoted by $B^{t}\colon W\to V$.
Following standard convention in the literature (see, e.g., [39]), we write $x\lesssim y$, or equivalently $y\gtrsim x$, to mean that there exists a constant $C>0$, independent of the relevant parameters, such that $x\leq Cy$. Furthermore, we write $x\eqsim y$ when both $x\lesssim y$ and $x\gtrsim y$ hold.
2 Auxiliary space theory
In this section, we provide a brief summary of the auxiliary space theory [33], which is a useful tool for the analysis of linear systems. In particular, we present sharp estimates for certain linear operators defined in terms of auxiliary spaces.
We begin with the auxiliary space lemma [33, Lemma 3.1], which is closely related to the fictitious space method [29] and the auxiliary space method [40]. This lemma plays an important role in the analysis developed in this paper. The idea of the lemma can be traced back to [39], and it has since been widely used in the design and analysis of iterative methods; see, for example, [1, 7].
Lemma 2.1 (auxiliary space lemma).
Let $V$ and $\undertilde{V}$ be finite-dimensional vector spaces and let $\Pi\colon\undertilde{V}\to V$ be a surjective linear operator. Let $\undertilde{B}\colon\undertilde{V}\to\undertilde{V}$ be a SPD linear operator, and define
(3) $B=\Pi\undertilde{B}\Pi^{t}\colon V\to V.$
Then $B$ is SPD. Moreover, it satisfies
$(B^{-1}v,v)=\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{B}^{-1}\undertilde{v},\undertilde{v}),\quad v\in V.$
Now, let $A\colon V\to V$ and $B\colon V\to V$ be SPD linear operators, and consider the situation where $B$ is used as a preconditioner for the linear system associated with AA. In this case, the convergence rate of iterative methods can be estimated in terms of the condition number of the preconditioned operator BABA, given by
$\kappa(BA)=\frac{\lambda_{\max}(BA)}{\lambda_{\min}(BA)},$
where $\lambda_{\min}$ and $\lambda_{\max}$ denote the minimum and maximum eigenvalues, respectively.
The following theorem, introduced in [33, Theorem 3.6] and proved using Lemma 2.1, provides sharp estimates for the extremal eigenvalues of BABA when the preconditioner $B$ is defined by (3) for some SPD linear operator $\undertilde{B}\colon\undertilde{V}\to\undertilde{V}$ and a surjective linear operator $\Pi\colon\undertilde{V}\to V$. Here, $\undertilde{V}$ is a finite-dimensional vector space, referred to as an auxiliary space. Thus, the preconditioner $B$ is constructed in terms of the operator $\undertilde{B}$ defined on the auxiliary space $\undertilde{V}$.
Theorem $2.$
Let $A\colon V\to V$ be a SPD linear operator, and let $B\colon V\to V$ be given in (3). Then we have
$\lambda_{\min}(BA)$ $=\left(\sup_{v\in V,\ \|v\|_{A}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{B}^{-1}\undertilde{v},\undertilde{v})\right)^{-1},$
$\lambda_{\max}(BA)$ $=\left(\inf_{v\in V,\ \|v\|_{A}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{B}^{-1}\undertilde{v},\undertilde{v})\right)^{-1}.$
The result in Theorem 2.2 can be extended to more general situations, for instance when AA is only semi-SPD, $B$ is a general linear operator, and $\undertilde{B}$ is semi-SPD. Further details on these extensions can be found in [33].
3 Saddle point problems
For completeness, we review the essentials of saddle point problems: the problem statement, well-posedness, the Schur complement system, and strategies for designing iterative methods. For comprehensive literature on saddle point problems, we refer the reader to, e.g., [2, 3].
3.1 Model problem
We consider the model saddle point problem (1). Since $B$ is surjective, we have (see [42])
(4) $\inf_{q\in W,\ \|q\|=1}\sup_{v\in V,\ \|v\|=1}(Bv,q)=\|B^{-1}\|^{-1}>0,$
where $B$ is regarded as an isomorphism from $\mathcal{N}(B)^{\perp}$ onto $W$. In other words, the Babuška–Brezzi (inf–sup) condition [3] holds.
Moreover, under the surjectivity assumption on $B$, the system operator
$\mathcal{A}=\begin{bmatrix}A&B^{t}\\ B&0\end{bmatrix}\colon V\times W\to V\times W$
associated with the problem (1) is nonsingular if and only if
(5) $\mathcal{N}(A)\cap\mathcal{N}(B)=\{0\},$
as established in [2, Theorem 3.4]. In what follows, we always assume that the condition (5) holds to ensure the well-posedness of (1).
3.2 Elimination of the primal variable
From the saddle point system (1), one can eliminate the primal variable $u$ and obtain a reduced system formulated solely in terms of the dual variable pp. Assume first that AA is SPD. From the first equation in (1), we have
(6) $u=A^{-1}(f-B^{t}p),$
and substituting (6) into the second equation of (1) yields the dual problem
(7) $Sp=d,\quad\text{where}\quad S=BA^{-1}B^{t}\colon W\to W,\quad d=BA^{-1}f-g\in W.$
In (7), $S$ is called the Schur complement. Because AA is SPD and $B$ is surjective, $S$ is of the form (3). In particular, by Lemma 2.1, we deduce that $S$ is SPD. Hence, the saddle point system (1) is equivalent to the SPD linear system (7) for $p$, followed by a back substitution to recover $u$.
The case when AA is only semi-SPD is more delicate, since (6) is not valid. Nevertheless, elimination of the primal variable $u$ is still possible when full information about the null space $\mathcal{N}(A)$ is available, that is, when an explicit basis of $\mathcal{N}(A)$ is known. Such a situation arises, for example, in Neumann boundary value problems for the Poisson equation and linear elasticity [11].
In the following, we summarize the arguments from [34, Section 2.2] and [37, Section 6.3] concerning the elimination of $u$. Let
$N\colon\mathbb{R}^{\dim\mathcal{N}(A)}\to\mathcal{N}(A)$
be an injective linear operator with range $\mathcal{N}(A)$. Such a map NN can be constructed as the matrix whose columns form a basis of $\mathcal{N}(A)$. From the first equation of (1), compatibility requires $f-B^{t}p\in\mathcal{R}(A)$, which is equivalent to
(8) $N^{t}(f-B^{t}p)=0.$
In this case, $u$ satisfies
(9) $u=A^{+}(f-B^{t}p)+N\xi,$
for some $\xi\in\mathbb{R}^{\dim\mathcal{N}(A)}$, where $A^{+}\colon V\to V$ denotes a pseudoinverse of AA, i.e., $AA^{+}$ acts as the identity on $\mathcal{R}(A)$. Substituting (9) into the second equation of (1) yields
(10) $Sp-BN\xi=d,\quad\text{where}\quad S=BA^{+}B^{t}\colon W\to W,\ d=BA^{+}f-g\in W.$
Note that (10) reduces to (7) when AA is SPD.
Now, (8) and (10) form a coupled system in the variables $p$ and $\xi$. We next eliminate $\xi$. Since $\mathcal{R}(N)=\mathcal{N}(A)$, the condition (5) implies that BNBN is injective. Hence the orthogonal projection onto
(11) $W_{0}:=\mathcal{R}(B|_{\mathcal{N}(A)})^{\perp}=\mathcal{R}(BN)^{\perp}\subset W$
is well defined and given by
(12) $P=I-BN((BN)^{t}BN)^{-1}(BN)^{t}\colon W\to W_{0}.$
The computation of $P$ requires solving a linear system of $(BN)^{t}BN$, which can be carried out at negligible cost when $\dim\mathcal{N}(A)$ is small.
We decompose $p$ orthogonally into components in $W_{0}$ and $W_{0}^{\perp}=\mathcal{R}(B|_{\mathcal{N}(A)})$:
(13) $p=P^{t}p_{0}+BN\eta,\quad p_{0}\in W_{0},\ \eta\in\mathbb{R}^{\dim\mathcal{N}(A)}.$
Substituting (13) into (8) gives
$\eta=((BN)^{t}BN)^{-1}N^{t}f.$
Note that $\eta$ can be computed prior to the main iterations of an iterative method. Finally, substituting (13) into (10) and applying $P$ to both sides yields
(14) $S_{0}p_{0}=d_{0},\quad\text{where}\quad S_{0}=PSP^{t}\colon W_{0}\to W_{0},\ d_{0}=d-PSBN\eta\in W_{0}.$
Thus, we arrive at a linear system in the single unknown $p_{0}$, which represents the dual problem in the case when AA is semi-SPD. Proposition 3.1 shows that $S_{0}$ is SPD and independent of the particular choice of the pseudoinverse $A^{+}$ (cf. [34, Lemma 2.19]).
Proposition 3.1.
The projected Schur complement $S_{0}\colon W_{0}\to W_{0}$ given in (14) is independent of the particular choice of the pseudoinverse $A^{+}$. Moreover, $S_{0}$ is SPD.
Proof 3.2.
Since $S_{0}=PBA^{+}(PB)^{t}$, to prove independence on the choice of $A^{+}$, it suffices to show that $PB|_{\mathcal{N}(A)}=0$, which follows directly from the definition (12) of $P$. Meanwhile, to prove that $S_{0}$ is SPD, it is enough, by Lemma 2.1, to verify that $PB|_{\mathcal{R}(A)}$ is surjective. This holds because PBPB is surjective and $PB|_{\mathcal{N}(A)}=0$.
Remark $3.$
While we define $P$ as the orthogonal projection onto $W_{0}$ in (12), it is also possible to employ oblique projections, which are often more effective for heterogeneous problems; see [34, 37]. However, we omit the details here.
Remark 3.4.
The dual problems (7) and (14) are special cases of Fenchel–Rockafellar duality in convex optimization, since (1) is equivalent to the quadratic optimization problem (2). For further details on Fenchel–Rockafellar duality, we refer the reader to [14].
3.3 Iterative methods
Here, we discuss iterative methods for solving the saddle point problem (1). For simplicity, we assume that AA is SPD. Based on the surveys [2, 41], there are three major approaches in iterative methods.
The first approach is that we apply iterative methods for the SPD dual problems (7) and (14). In this case, we can utilize vast existing results on iterative methods for SPD linear systems. See [33] for a systematic approach for design and analysis of iterative methods for SPD linear systems.
The second approach is to apply stationary iterations, such as Uzawa-type methods [4] or the augmented Lagrangian method [12], to (1). It is well-known that the classical Uzawa method and the augmented Lagrangian method are equivalent to certain iterative schemes for solving the dual problem (7[2]. Moreover, for general inexact Uzawa methods, their convergence behavior is determined by properties of the Schur complement [4].
The third approach is to apply preconditioned Krylov subspace methods, such as MINRES [31], to (1). It was shown in [28] that the block-diagonal preconditioner
$\mathcal{L}=\begin{bmatrix}A^{-1}&0\\ 0&-S^{-1}\end{bmatrix}$
is optimal in the sense that the preconditioned operator $\mathcal{L}\mathcal{A}$ has only three distinct eigenvalues, namely 11 and $\tfrac{1}{2}(1\pm\sqrt{5})$. Thus, in this case, the central task is the construction of efficient preconditioners for AA and $S$ [2].
In all three approaches, the properties of the Schur complement $S$ are central to both design and analysis of algorithms. Hence, in the remainder of this paper, we focus on the analysis of $S$. In particular, we investigate how $S$ can be studied in a concise and sharp manner using the auxiliary space theory introduced in Section 2.
4 Estimates for the Schur complements
In this section, we present useful estimates for the extremal eigenvalues of the Schur complements given in (7) and (14). Although these estimates are derived in a relatively straightforward manner using the auxiliary space theory, they provide bounds that are equivalent to, or even sharper than, existing results, for which more complicated proofs were often required.
We begin with the case when AA is SPD. An obvious but important observation is that the Schur complement $S$ has the same structure as in (3). In Theorem 4.1, we establish sharp estimates for the extremal eigenvalues of $S$ defined in (7), obtained as direct consequences of Theorem $2.$
Theorem 4.1.
In the saddle point problem (1), suppose that AA is SPD. Then the Schur complement $S$ given in (7) satisfies
$\lambda_{\min}(S)=\inf_{0\neq q\in W}\sup_{v\in V,\ Bv=q}\frac{\|q\|^{2}}{(Av,v)},\quad\lambda_{\max}(S)=\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{(Av,v)}.$
Proof 4.2.
In Theorem 2.2, we set
$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad A\leftarrow I,\quad B\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{B}\leftarrow A^{-1}.$
Then we get
$\lambda_{\min}(S)=\left(\sup_{0\neq q\in W}\inf_{v\in V,\ Bv=q}\frac{(Av,v)}{\|q\|^{2}}\right)^{-1}=\inf_{0\neq q\in W}\sup_{v\in V,\ Bv=q}\frac{\|q\|^{2}}{(Av,v)}.$
Moreover, we have
$\lambda_{\max}(S)=\left(\inf_{0\neq q\in W}\inf_{v\in V,\ Bv=q}\frac{(Av,v)}{\|q\|^{2}}\right)^{-1}=\left(\inf_{0\neq v\in V}\frac{(Av,v)}{\|Bv\|^{2}}\right)^{-1}=\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{(Av,v)},$
which completes the proof.
As a corollary of Theorem 4.1, we obtain Corollary 4.3, which appeared in [37, Lemma 9.1] (see also [3]).
Corollary 4.3.
In the saddle point problem (1), suppose that AA is SPD. Then the Schur complement $S$ given in (7) satisfies
$\lambda_{\min}(S)\geq\lambda_{\max}(A)^{-1}\|B^{-1}\|^{-2},\quad\lambda_{\max}(S)\leq\lambda_{\min}(A)^{-1}\|B\|^{2}.$
Proof $4.$
By Theorem 4.1, we obtain
$\lambda_{\min}(S)=\inf_{0\neq q\in W}\sup_{v\in V,\ Bv=q}\frac{\|q\|^{2}}{(Av,v)}\geq\lambda_{\max}(A)^{-1}\inf_{0\neq q\in W}\sup_{v\in V,\ Bv=q}\frac{\|q\|^{2}}{\|v\|^{2}}\\ =\lambda_{\max}(A)^{-1}\left(\sup_{0\neq q\in W}\inf_{v\in V,\ Bv=q}\frac{\|v\|^{2}}{\|q\|^{2}}\right)^{-1}=\lambda_{\max}(A)^{-1}\|B^{-1}\|^{-2},$
where in the last inequality we use the fact that the minimizer $v\in V$ of $\|v\|^{2}/\|q\|^{2}$ subject to $Bv=q$ must lie in $\mathcal{N}(B)^{\perp}$. Meanwhile, we have
$\lambda_{\max}(S)=\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{(Av,v)}\leq\lambda_{\min}(A)^{-1}\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{\|v\|^{2}}=\lambda_{\min}(A)^{-1}\|B\|^{2},$
which completes the proof.
Sometimes a preconditioner $L\colon W\to W$ is applied to the Schur complement $S$. Theorem 4.5 provides sharp estimates for the extremal eigenvalues of the preconditioned operator LSLS.
Theorem 4.5.
In the saddle point problem (1), suppose that AA is SPD. Let $L\colon W\to W$ be a SPD linear operator. Then the Schur complement $S$ given in (7) satisfies
$\lambda_{\min}(LS)=\inf_{0\neq q\in W}\sup_{v\in V,\ Bv=q}\frac{(Lq,q)}{(Av,v)},\quad\lambda_{\max}(LS)=\sup_{0\neq v\in V}\frac{(LBv,Bv)}{(Av,v)}.$
Proof 4.6.
Setting
$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad A\leftarrow L,\quad B\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{B}\leftarrow A^{-1}$
in Theorem 2.2 yields the desired result.
An interesting special case arises when a right inverse of $B$ is explicitly available. In this situation, one can construct a special preconditioner $L$ using the right inverse of $B$ as a building block; examples include certain nonoverlapping domain decomposition methods, which we will discuss in Section 7. The following corollary, originally stated in [25, Lemma 2] with a rather involved proof, follows directly from Theorem 4.5.
Corollary 4.7.
In the saddle point problem (1), suppose that AA is SPD. Assume further that a linear operator $\bar{B}\colon V\to W$ satisfies $B\bar{B}^{t}=I$, and define a preconditioner $L\colon W\to W$ by
$L=\bar{B}A\bar{B}^{t}.$
Then the Schur complement $S$ given in (7) satisfies
$\lambda_{\min}(LS)\geq 1,\quad\lambda_{\max}(LS)=\sup_{0\neq v\in V}\frac{(A\bar{B}^{t}Bv,\bar{B}^{t}Bv)}{(Av,v)}=\|\bar{B}^{t}B\|_{A}^{2},$
where $\|\cdot\|_{A}$ denotes the operator norm with respect to the AA-norm on $V$.
Proof 4.8.
The estimate for the minimum eigenvalue follows by substituting $v=\bar{B}^{t}q$ into the expression in Theorem 4.5, while the estimate for the maximum eigenvalue is an immediate consequence of Theorem 4.5.
Next, we consider the case when AA is semi-SPD. Recall that the projected Schur complement $S_{0}$ given in (14) satisfies
$S_{0}=(PB)A^{+}(PB)^{t}.$
As discussed in the proof of Proposition 3.1, the operator $PB|_{\mathcal{R}(A)}$ is surjective and $\mathcal{R}((PB)^{t})\subset\mathcal{R}(A)$. Hence, in view of Proposition 3.1, we may regard $A^{+}$ and PBPB as
$A^{+}\colon\mathcal{R}(A)\to\mathcal{R}(A),\quad PB\colon\mathcal{R}(A)\to W_{0},$
where $W_{0}$ was defined in (11). This observation shows that the analysis of the semi-SPD case proceeds in essentially the same way as in the SPD case; see Theorems 4.9 and 4.10.
Theorem 4.9.
In the saddle point problem (1), suppose that AA is semi-SPD. Then the projected Schur complement $S_{0}$ given in (14) satisfies
$\lambda_{\min}(S_{0})=\inf_{0\neq q\in W_{0}}\sup_{v\in\mathcal{R}(A),\ PBv=q}\frac{\|q\|^{2}}{(Av,v)},\quad\lambda_{\max}(S_{0})=\sup_{0\neq v\in\mathcal{R}(A)}\frac{\|PBv\|^{2}}{(Av,v)}.$
Theorem 4.10.
In the saddle point problem (1), suppose that AA is semi-SPD. Let $L\colon W_{0}\to W_{0}$ be a SPD linear operator. Then the projected Schur complement $S_{0}$ given in (14) satisfies
$\lambda_{\min}(LS_{0})=\inf_{0\neq q\in W_{0}}\sup_{v\in\mathcal{R}(A),\ PBv=q}\frac{(Lq,q)}{(Av,v)},\quad\lambda_{\max}(LS_{0})=\sup_{0\neq v\in\mathcal{R}(A)}\frac{(LPBv,PBv)}{(Av,v)}.$
Proof 4.11.
Setting
$V\leftarrow W_{0},\quad\undertilde{V}\leftarrow\mathcal{R}(A),\quad A\leftarrow L,\quad B\leftarrow S_{0},\quad\Pi\leftarrow PB,\quad\undertilde{B}\leftarrow A^{+}$
in Theorem 2.2 yields the desired result.
For completeness, we also provide a counterpart of Corollary 4.7 for the semi-SPD case in Corollary 4.12.
Corollary 4.12.
In the saddle point problem (1), suppose that AA is semi-SPD. Assume further that a linear operator $\bar{B}\colon V\to W$ satisfies $B\bar{B}^{t}=I$, and define a preconditioner $L\colon W_{0}\to W_{0}$ by
$L=(P\bar{B})A(P\bar{B})^{t}.$
Then the projected Schur complement $S_{0}$ given in (14) satisfies
$\lambda_{\min}(LS_{0})\geq 1,\quad\lambda_{\max}(LS_{0})=\sup_{0\neq v\in\mathcal{R}(A)}\frac{(A(P\bar{B})^{t}P\bar{B}v,(P\bar{B})^{t}P\bar{B}v)}{(Av,v)}.$
5 Application I: Augmented Lagrangian method
In the remainder of this paper, we present several examples illustrating how the auxiliary space theory and the estimates for the Schur complements developed in Section 4 can be applied to different applications. As a first example, we consider the augmented Lagrangian method [12], which is a stationary iterative method for solving the saddle point problem (1).
In the augmented Lagrangian method, given the current iterate $(u^{m},p^{m})$, the next iterate $(u^{m+1},p^{m+1})$ is obtained as
(15) $\begin{aligned} u^{m+1}&=(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon B^{t}g-B^{t}p^{m}),\\ p^{m+1}&=p^{m}-\epsilon^{-1}(g-Bu^{m+1}),\end{aligned}\quad m\geq 0.$
It is well-known (see, e.g., [2]) that the augmented Lagrangian method (15) is equivalent to the Richardson iteration with step size $\epsilon^{-1}$ applied to the linear system
(16) $S_{\epsilon}p=d_{\epsilon},\quad\text{where}\quad S_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}B^{t},\,d_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon B^{t}g)-g.$
Therefore, the analysis of the augmented Lagrangian method reduces to the analysis of the augmented Schur complement $S_{\epsilon}$.
In the following, we derive some identities for $S_{\epsilon}$, which can be regarded as refinements of existing results [18, 21], by invoking the auxiliary space theory. By repeated applications of Lemma 2.1, for any $q\in W$ we have
$(S_{\epsilon}^{-1}q,q)=\inf_{v\in V,\ Bv=q}((A+\epsilon^{-1}B^{t}B)v,v)\\ =\inf_{v\in V,\ Bv=q}(Av,v)+\epsilon^{-1}(q,q)=(S^{-1}q,q)+\epsilon^{-1}(q,q).$
Thus we obtain the identity (cf. [18, equation (3.4)]):
$S_{\epsilon}^{-1}=S^{-1}+\epsilon^{-1}I.$
Consequently, the extremal eigenvalues of $S_{\epsilon}$ are
(17) $\lambda_{\min}(S_{\epsilon})=\frac{\epsilon\lambda_{\min}(S)}{\epsilon+\lambda_{\min}(S)},\quad\lambda_{\max}(S_{\epsilon})=\frac{\epsilon\lambda_{\max}(S)}{\epsilon+\lambda_{\max}(S)}.$
In the augmented Lagrangian method (15), the error propagation operator is given by $I-\epsilon^{-1}S_{\epsilon}$. Using (17), we have the following sharp estimate:
$\|I-\epsilon^{-1}S_{\epsilon}\|=\lambda_{\max}(I-\epsilon^{-1}S_{\epsilon})=1-\epsilon^{-1}\lambda_{\min}(S_{\epsilon})=\frac{\epsilon}{\epsilon+\lambda_{\min}(S)}.$
Namely, the convergence rate of (15) becomes arbitrarily fast as $\epsilon$ becomes small. This result can be regarded as a refined version of [21, Lemma 2.1], with much simpler proof. We note that such arbitrarily fast convergence of the augmented Lagrangian method can be extended to a broader class of convex optimization problems; see [32, Appendix A].
On the other hand, if we apply the conjugate gradient method to solve (16), then the convergence rate depends on the condition number $\kappa(S_{\epsilon})$, which satisfies
$\kappa(S_{\epsilon})=\frac{\lambda_{\max}(S)}{\lambda_{\min}(S)}\frac{\epsilon+\lambda_{\min}(S)}{\epsilon+\lambda_{\max}(S)}\rightarrow 1\quad\text{as}\quad\epsilon\to 0.$
This result both refines and agrees with [17, 18], where the condition number of the augmented Lagrangian method applied to nonoverlapping domain decomposition methods was analyzed.
6 Application II: Mixed finite element methods
In this section, we present an analysis of several mixed finite element methods for solving mixed variational problems [3]. As representative examples, we consider Darcy flow [5, 32] and the Stokes equations [1, 38], which have broad applications in fluid mechanics.
6.1 Darcy flow
Let $\Omega$ be a bounded polyhedral domain in $\mathbb{R}^{d}$ ($d=2,3$). The linear relationship between the Darcy velocity $u$ and pressure $p$, together with the conservation of mass and the homogeneous Dirichlet boundary condition, is given by
(18) $u+\nabla p$ $=0\quad\text{in }\Omega,$
$\operatorname{div}u$ $=b\quad\text{in }\Omega,$
$p$ $=0\quad\text{on }\partial\Omega,$
where $b$ is a prescribed source or sink term. Problems of the same form also arise in the mixed formulation of the Poisson equation, as well as in the Hellinger–Reissner variational principle for linear elasticity.
The weak formulation of (18) defined on $H(\operatorname{div};\Omega)\times L^{2}(\Omega)$ reads as follows: find $u\in H(\operatorname{div};\Omega)$ and $p\in L^{2}(\Omega)$ such that
(19) $\begin{aligned} \int_{\Omega}u\cdot v\,dx-\int_{\Omega}p\,\operatorname{div}v\,dx&=0,\\ -\int_{\Omega}q\,\operatorname{div}u\,dx&=-\int_{\Omega}bq\,dx,\end{aligned}\quad v\in H(\operatorname{div};\Omega),\ q\in L^{2}(\Omega).$
A mixed finite element method for solving (19) is obtained by replacing $H(\operatorname{div};\Omega)$ and $L^{2}(\Omega)$ with suitable finite element spaces $V$ and $W$. The corresponding algebraic system is
(20) $\begin{bmatrix}M_{V}&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ p\end{bmatrix}=\begin{bmatrix}0\\ g\end{bmatrix}$,
where the matrices $M_{V}$, $B$, and the vector $g$ are defined by
$(M_{V}v,w)=\int_{\Omega}v\cdot w\,dx,\ (Bv,q)=-\int_{\Omega}q\,\operatorname{div}v\,dx,\ (g,q)=-\int_{\Omega}bq\,dx,\quad v,w\in V,\ q\in W,$
and $(\cdot,\cdot)$ denotes the Euclidean inner product. Thus, the problem (20) is an instance of (1).
It is straightforward to verify that
(21) $B=-M_{W}\operatorname{div},$
where $M_{W}\colon W\to W$ is the mass matrix on $W$ defined analogously to $M_{V}$. Moreover, the mass matrices $M_{V}$ and $M_{W}$ satisfy the scaling relations (cf. [37, Section 3.4])
(22) $(M_{V}v,v)\eqsim h^{d}\|v\|^{2},\quad(M_{W}q,q)\eqsim h^{d}\|q\|^{2},\quad v\in V,\ q\in W,$
where $h>0$ denotes the characteristic element diameter.
We assume that the finite element spaces $V$ and $W$ satisfy the discrete Babuška–Brezzi condition (see, e.g., [5]), namely,
$\inf_{q\in W}\sup_{v\in V}\frac{(\operatorname{div}v,q)_{L^{2}}}{\|v\|_{H(\operatorname{div})}\|q\|_{L^{2}}}\gtrsim 1.$
By (4), this condition is equivalent to the statement that for any $q\in W$, there exists $v\in V$ such that
(23) $\operatorname{div}v=q,\quad\|q\|_{L^{2}}\gtrsim\|v\|_{H(\operatorname{div})}.$
Now, we analyze the Schur complement $S=BM_{V}^{-1}B^{t}$ corresponding to (20). By Theorem 4.1, the minimum eigenvalue of $S$ is estimated as follows:
[…]
For the maximum eigenvalue, we have
[…]
where the last inequality follows from the standard inverse inequality (see, e.g., [30, Lemma 2.2]). In conclusion, we obtain
$\kappa(S)\lesssim h^{-2}.$
Since $\kappa(S)$ grows as the element diameter $h$ decreases, a preconditioner is required. As discussed in [41], an optimal preconditioner for $S$ can be constructed by exploiting its spectral equivalence with a certain discretization of the Poisson problem. Indeed, a preconditioner based on the spectral equivalence with the Poisson problem discretized by an interior penalty method was considered in [35].
6.2 Stokes equations
Let $\Omega$ be a bounded polygonal domain in $\mathbb{R}^{d}$ ($d=2,3$). The incompressible Stokes equations with the homogeneous Dirichlet boundary condition are written as
(24) $-\Delta u+\nabla p$ $=f\quad$ $\text{in }\Omega,$
$\operatorname{div}u$ $=0\quad$ $\text{in }\Omega,$
$u$ $=0\quad$ $\text{on }\partial\Omega,$
where $u$ denotes the velocity, $p$ the pressure, and $f$ the external body force density.
The weak formulation of (24) on $H_{0}^{1}(\Omega)^{d}$ and $L_{0}^{2}(\Omega)$ is given as follows: find $(u,p)\in H_{0}^{1}(\Omega)^{d}\times L_{0}^{2}(\Omega)$ such that
(25) $\begin{aligned} \int_{\Omega}\nabla u\cdot\nabla v\,dx-\int_{\Omega}p\,\operatorname{div}v\,dx&=\int_{\Omega}f\cdot v\,dx,\\ -\int_{\Omega}q\,\operatorname{div}u\,dx&=0,\end{aligned}\quad v\in H_{0}^{1}(\Omega)^{d},\ q\in L_{0}^{2}(\Omega).$
A mixed finite element method for solving (24) is obtained by replacing $H_{0}^{1}(\Omega)^{d}$ and $L_{0}^{2}(\Omega)$ with suitable finite element spaces $V$ and $W$. Then the corresponding algebraic system becomes (1), where the matrices AA, $B$ and the vector $f$ are defined by
$(Av,w)=\int_{\Omega}\nabla v\cdot\nabla w\,dx,\ (Bv,q)=-\int_{\Omega}q\operatorname{div}v\,dx,\ (f,v)=\int_{\Omega}f\cdot v\,dx,\quad v,w\in V,\ q\in W.$
Let $M_{W}\colon W\to W$ be the mass matrix on $W$. Then, similarly to the case of Darcy flow, we observe that (21) and (22) hold. Moreover, if we assume that the finite element spaces $V$ and $W$ satisfy the discrete Babuška–Brezzi condition (see [3]),
$\inf_{q\in W}\sup_{v\in V}\frac{(\operatorname{div}v,q)_{L^{2}}}{\|v\|_{H^{1}}\|q\|_{L^{2}}}\gtrsim 1,$
then by (4), this condition is equivalent to the statement that for any $q\in W$, there exists $v\in V$ such that
(27) $\operatorname{div}v=q,\quad\|q\|_{L^{2}}\gtrsim\|v\|_{H^{1}}.$
Now, we estimate the extremal eigenvalues of the Schur complement $S=BA^{-1}B^{t}$. By Theorem 4.1, the minimum eigenvalue of $S$ is estimated as follows:
[…]
On the other hand, the maximum eigenvalue of $S$ is estimated as follows:
[…]
Consequently, we obtain
$\kappa(S)\lesssim 1.$
This implies that the conjugate gradient method for solving the dual problem converges uniformly with respect to the characteristic element diameter $h$, which agrees with [38].
7 Application III: FETI and FETI-DP
In this section, we present applications of the proposed framework to two widely used nonoverlapping domain decomposition methods [37, 43]: FETI [11] and FETI-DP [9, 16]. Nonoverlapping domain decomposition methods have been actively studied for several decades. Although these methods are highly efficient in practice, their mathematical analysis is often rather involved. We show how the abstract estimates for the Schur complements developed in Section 4 can lead to much simpler analyses of these methods.
7.1 FETI
We first consider FETI, originally proposed in [11]. In FETI, a global problem is partitioned into smaller subproblems posed on subdomains, and continuity across subdomain interfaces is enforced by means of Lagrange multipliers.
To illustrate, let $\Omega$ be a bounded polygonal domain in $\mathbb{R}^{2}$, and consider the model Poisson equation
(28) $-\Delta u$ $=f\quad\text{in }\Omega,$
$u$ $=0\quad\text{on }\partial\Omega.$
Assume that the domain $\Omega$ is decomposed into $J$ nonoverlapping polygonal subdomains $\{\Omega_{j}\}_{j=1}^{J}$ with characteristic subdomain diameter $H>0$, where each $\Omega_{j}$ is a union of coarse elements from a global conforming mesh $\mathcal{T}_{H}$. We further assume that each coarse element in $\mathcal{T}_{H}$ is itself a union of fine elements from a fine mesh $\mathcal{T}_{h}$. For $i<j$, let $\Gamma_{ij}=\partial\Omega_{i}\cap\partial\Omega_{j}$ denote the interface between adjacent subdomains, and define $\Gamma=\bigcup_{i<j}\Gamma_{ij}$.
In what follows, the indices $i<j$ are taken over $1,\dots,J$. Let $\mathcal{T}_{h}^{j}$ denote the restriction of the global triangulation $\mathcal{T}_{h}$ to $\Omega_{j}$. Then $\mathcal{T}_{h}^{i}$ and $\mathcal{T}_{h}^{j}$ share nodal points along the interface $\Gamma_{ij}$. On each $\Omega_{j}$, we consider the space of continuous, piecewise linear finite elements on $\mathcal{T}_{h}^{j}$ that vanish on $\partial\Omega_{j}\cap\partial\Omega$, and denote its restriction to the interface $\partial\Omega_{j}$ by $V_{j}$. We also define the product space $V=\prod_{j=1}^{J}V_{j}$. Note that functions in $V$ are, in general, discontinuous across $\Gamma$.
The problem of interest is
(29) $\min_{v=(v_{j})_{j=1}^{J}\in V}\left\{\frac{1}{2}(Sv,v)-(f,v)\right\}\quad\text{subject to}\quad Bv=0,$
where the matrix $S$ and the vector $f$ are defined by
(30) $(Sv,w)=\sum_{j=1}^{J}\int_{\Omega_{j}}\nabla\mathcal{H}_{j}v_{j}\cdot\nabla\mathcal{H}_{j}w_{j}\,dx,\qquad(f,v)=\sum_{j=1}^{J}\int_{\Omega_{j}}f\,\mathcal{H}_{j}v_{j}\,dx,\quad v,w\in V,$
$\mathcal{H}_{j}$ denotes the discrete harmonic extension in $\Omega_{j}$ associated with $\mathcal{T}_{h}^{j}$ (see [37, 43] for details), $B$ is a full-rank matrix with entries 0 and $\pm 1$ enforcing continuity along $\Gamma$, and $(\cdot,\cdot)$ denotes the Euclidean inner product. It is straightforward to observe that (29) is equivalent to the restriction of the finite element discretization of (28) on the global mesh $\mathcal{T}_{h}$ to the interface $\Gamma$.
By enforcing the constraint $Bv=0$ in (29) through the introduction of a Lagrange multiplier $\lambda\in W$, where $W$ denotes the vector space of Lagrange multipliers, we obtain the following saddle point problem, which is of the form (1):
(31) $\begin{bmatrix}S&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}f\\ 0\end{bmatrix}$.
Note that $S$ is semi-SPD, owing to subdomains $\Omega_{j}$ that do not intersect $\partial\Omega$. Finally, by eliminating $u$ in (31) as in Section 4, we obtain the following SPD linear system corresponding to (14):
(32) $F_{0}\lambda_{0}=d_{0},\quad\text{where}\quad F_{0}=PBS^{+}(PB)^{t}\colon W_{0}\to W_{0},\ d_{0}\in W_{0},$
and $W_{0}$ and $P$ are defined in the same way as in (11) and (12), respectively.
Now we estimate the condition number of $F_{0}$ in (32) using the proposed framework. We begin by summarizing some useful properties of the matrices $S$ and $B$. The matrix $S$ satisfies the following bounds, which combine the well-known estimate for Schur complements in the finite element method (see, e.g., [34, Lemma 1.59]) with standard scaling arguments [37, Section 3.4]:
(33) $\lambda_{\min}(S)\gtrsim\frac{h}{H},\quad\lambda_{\max}(S)\lesssim 1,$
where $\lambda_{\min}$ denotes the minimum nonzero eigenvalue. Furthermore, as shown in [10, Equation (22)], since the matrix $B$ consists only of entries 0 and $\pm 1$, an elementary calculation gives
(34) $\lambda_{\min}(BB^{t})\gtrsim 1,\quad\lambda_{\max}(BB^{t})=\lambda_{\max}(B^{t}B)\lesssim 1.$
By Theorem 4.9, the minimum eigenvalue of $F_{0}$ is estimated as
[…]
where in the second inequality we substituted $v=B^{t}(BB^{t})^{-1}\mu$, which satisfies $PBv=\mu$. On the other hand, the maximum eigenvalue of $F_{0}$ is estimated as
[…]
Consequently, we get
$\kappa(F_{0})\lesssim\frac{H}{h}.$
This result agrees with [10, Theorem 3.2], where the condition number of $F_{0}$ was first estimated.
Next, we consider the following scaled Dirichlet preconditioner:
$L_{\mathrm{FETI}}=PB_{D}S(PB_{D})^{t},$
where $B_{D}=D_{\mathrm{FETI}}B$ is a scaled jump matrix [36, 37], with $D_{\mathrm{FETI}}$ a block-diagonal scaling matrix chosen such that $BB_{D}^{t}=I$. In other words, $B_{D}^{t}$ is a right inverse of $B$. In this case, we can apply Corollary 4.12 to estimate the extremal eigenvalues of the preconditioned operator $L_{\mathrm{FETI}}S_{0}$.
The only technical ingredient we require is the well-known vertex–edge lemma [43, Section 4.3], stated in Lemma 7.1. In what follows, edges are understood as relatively open subsets, i.e., without their endpoints. For a subset $K\subset\partial\Omega_{j}$, the restriction $I_{K}^{0}v_{j}\in V_{j}$ of $v_{j}\in V_{j}$ is defined by
[…]
Lemma 7.1.
For each subdomain $\Omega_{j}\subset\mathbb{R}^{2}$, the following estimates hold:
  1. (a)
    If $x$ is a vertex of $\partial\Omega_{j}$, then we have
    $\|I_{x}^{0}v\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)\|v\|_{H^{\frac{1}{2}}(\partial\Omega)}^{2},\quad v\in V_{j}.$
  2. (b)
    If $e$ is an edge of $\partial\Omega_{j}$, then we have
    $\|I_{e}^{0}v\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)^{2}\|v\|_{H^{\frac{1}{2}}(\partial\Omega)}^{2},\quad v\in V_{j}.$
Now we are ready to analyze the preconditioned operator $L_{\mathrm{FETI}}S_{0}$. By Corollary 4.12, we have
$\lambda_{\min}(L_{\mathrm{FETI}}S_{0})\geq 1.$
To estimate the maximum eigenvalue of $L_{\mathrm{FETI}}S_{0}$, Corollary 4.12 implies that it suffices to bound $(S(P\bar{B})^{t}PBv,(P\bar{B})^{t}PBv)$ for any $v\in\mathcal{R}(S)$. Take any $v\in\mathcal{R}(S)$ and choose $\phi\in\mathcal{N}(S)$ such that $v+\phi\in W_{0}$. Then we have
(36) $P^{t}PB(v+\phi)=B(v+\phi).$
It follows that
[…]
where the first inequality follows from the definitions of $B$ and $\bar{B}$ together with the triangle inequality, and the second inequality follows from Lemma 7.1. Here we require mild geometric assumptions on the domain decomposition $\{\Omega_{j}\}_{j=1}^{J}$, namely that the numbers of vertices and edges of the subdomains are uniformly bounded and that the number of subdomains sharing a common vertex is also uniformly bounded. Hence, by Corollary 4.12, we deduce that
$\lambda_{\max}(L_{\mathrm{FETI}}F_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
In summary, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{FETI}}F_{0}$ (cf. [37, Theorem 6.15]):
$\kappa(L_{\mathrm{FETI}}F_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
We note that the unscaled Dirichlet preconditioner, namely $(PB)S(PB)^{t}$, was first proposed in [10] and analyzed in [26]. In the unscaled case, as discussed in [26], the minimum eigenvalue admits a weaker estimate in certain scenarios.
7.2 FETI-DP
Next, we consider FETI-DP [9, 16], which is one of the most widely used variants of FETI. It has been applied not only to linear problems but also to more complex nonlinear problems arising in computational mechanics [15, 20].
While the FETI formulation (29) enforces continuity along the entire interface $\Gamma$ through the constraint, FETI-DP instead imposes continuity at subdomain corners directly by restricting the solution space to a subspace $\tilde{V}\subset V$. The space $\tilde{V}$ consists of functions in $V$ that are continuous at subdomain corners. Continuity along the interior of each subdomain edge is enforced by constraints. Namely, we consider the following formulation:
(37) $\min_{v=(v_{j})_{j=1}^{J}\in\tilde{V}}\left\{\frac{1}{2}(\tilde{S}v,v)-(f,v)\right\},\quad\text{subject to}\quad Bv=0,$
where the matrix $\tilde{S}$ and the vector $\tilde{f}$ are defined analogously to (30); see [16, 37] for details. Here $B$ is a full-rank matrix with entries 0 and $\pm 1$ that enforces continuity along the interior of subdomain edges, and $(\cdot,\cdot)$ denotes the Euclidean inner product.
By enforcing the constraint $Bv=0$ in (37) through the introduction of a Lagrange multiplier $\lambda\in W$, where $W$ denotes the vector space of Lagrange multipliers, we obtain the saddle point problem
(38) $\begin{bmatrix}\tilde{S}&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}f\\ 0\end{bmatrix}$.
Different from the FETI formulation (31), in (38) the matrix $\tilde{S}$ is SPD. Therefore, by eliminating $u$ in (38), we obtain the following Schur complement system corresponding to (7):
(39) $F\lambda=d,\quad\text{where}\quad F=B\tilde{S}^{-1}B^{t},\quad d=\tilde{S}^{-1}f.$
In (37), since the matrix $B$ enforces continuity only at nodal points located in the interior of subdomain edges, each nodal point is affected by $B$ only once. Hence, we have the following identity [27, Equation (3.2)]:
(40) $BB^{t}=2I.$
Thus, if we define $\bar{B}=\tfrac{1}{2}B$, then $\bar{B}^{t}$ is a right inverse of $B$.
As FETI-DP involves functions that are continuous at subdomain corners, we require certain Poincaré-type inequalities associated with subdomain corners in order to analyze FETI-DP. In Lemma 7.2, which summarizes results from [19, Lemma 2] and [27, Lemma 5.1], $L^{2}$ and $H^{\frac{1}{2}}$ estimates are provided for functions vanishing at subdomain corners. In what follows, let $I_{H}$ denote the linear coarse interpolation operator associated with the coarse triangulation $\mathcal{T}_{H}$.
Lemma 7.2.
For each subdomain $\Omega_{j}\subset\mathbb{R}^{2}$, the following estimates hold:
$\begin{aligned} \|v_{j}-I_{H}v_{j}\|_{L^{2}(\partial\Omega_{j})}^{2}\lesssim H\left(1+\log\frac{H}{h}\right)|v_{j}|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2},\\ \sum_{e\text{:~edge of }\Omega_{j}}|I_{e}^{0}(v_{j}-I_{H}v_{j})|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)^{2}|v_{j}|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2},\end{aligned}\quad v_{j}\in V_{j},$
where $I_{e}^{0}$ was defined in (35).
Now, we estimate the extremal eigenvalues of $F$ given in (39). By Theorem 4.1, we obtain
[…]
On the other hand, since $\mathcal{R}(I_{H})\subset\mathcal{N}(B)$, for any $v=(v_{j})_{j=1}^{J}\in\tilde{V}$, we have
[…]
where the last inequality is due to Lemma 7.2. Hence, Theorem 4.1 implies that
$\lambda_{\max}(F)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right).$
Consequently, we deduce that
$\kappa(F)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right),$
which agrees with the existing result in [19, Proposition 3].
Next, we consider the following Dirichlet preconditioner:
$L_{\mathrm{DP}}=\bar{B}\tilde{S}\bar{B}^{t},$
By Corollary 4.7, we have $\lambda_{\min}(L_{\mathrm{DP}}F)\geq 1$. Moreover, for any $v=(v_{j})_{j=1}^{J}\in\tilde{V}$, it follows that
$\begin{split}(\tilde{S}\bar{B}^{t}Bv,\bar{B}^{t}Bv)&=(\tilde{S}\bar{B}^{t}B(v-I_{H}v),\bar{B}^{t}B(v-I_{H}v))\\ &=\sum_{j=1}^{J}|\bar{B}^{t}B(v-I_{H}v)|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\sum_{j=1}^{J}\sum_{e\text{:~edge of }\Omega_{j}}\left|I_{e}^{0}(v_{j}-I_{H}v_{j})\right|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\left(1+\log\frac{H}{h}\right)^{2}\sum_{j=1}^{J}|v_{j}|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &=\left(1+\log\frac{H}{h}\right)^{2}(\tilde{S}v,v),\end{split}$
where the first inequality follows from the definitions of $B$ and $\bar{B}$ together with the triangle inequality, and the second inequality is due to Lemma 7.2. This implies
$\lambda_{\max}(L_{\mathrm{DP}}F)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
Finally, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{DP}}F$ (cf. [27, Theorem 5.2]):
$\kappa(L_{\mathrm{DP}}F)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
The above proof can be extended to the cases of three dimensions or heterogeneous coefficients [16, 37] without major difficulties.
Remark 7.3.
Two closely related nonoverlapping domain decomposition methods to FETI and FETI-DP are BDD (balancing domain decomposition) [13, 23] and BDDC (balancing domain decomposition by constraints) [8, 13]. It has been shown in the literature [6, 22, 24, 25, 36] that FETI/FETI-DP with scaled Dirichlet preconditioners, and BDD/BDDC are spectrally equivalent in the sense that their spectra agree except for the eigenvalues 0 and 11. In fact, BDD and BDDC can be analyzed within our framework (Theorems 4.5 and 4.10) by essentially the same arguments used for FETI and FETI-DP. For brevity, we omit the details.
8 Concluding remarks
In this paper, we have presented sharp estimates for the Schur complements arising in saddle point problems, by utilizing the auxiliary space theory introduced in [33]. We also provided various examples, such as the augmented Lagrangian method, mixed finite element methods, and nonoverlapping domain decomposition methods, which highlight the usefulness of our results. In particular, the proposed framework simplifies the analysis compared to existing approaches by minimizing the algebraic effort required. We believe that the framework not only offers deeper insight into saddle point problems but also serves as a valuable tool for the design and analysis of new efficient iterative methods.
We conclude the paper with two remarks. The first concerns the case when the operator $B$ in (1) is not surjective, so that (1) becomes singular and admits nonunique solutions. Such a situation arises, for example, when nonredundant Lagrange multipliers are used to enforce the constraint $Bv=g$ in (2) (see, e.g., [34, 37]). In this case, the Schur complement $S$ becomes semi-SPD. Although we do not treat this case in detail here for the sake of simplicity and readability, we note that it can be handled by essentially the same arguments presented in this paper, since the auxiliary space theory [33] provides a unified analysis of both SPD and semi-SPD problems.
The second remark concerns the extension of the proposed framework to a broader range of problems. As mentioned at several points in this paper (see, e.g., Remark 3.4), some parts of the contents of this paper can already be extended to more general classes of convex optimization problems, although a full generalization is highly nontrivial. We note that there has recently been significant research on extending the theory of iterative methods for linear systems to convex optimization problems; see [14] and the references therein. In this spirit, we view the extension of our results to convex optimization problems as an important direction for future work, which we expect to be highly impactful.

### v2

@@
 The purpose of this paper is to present sharp estimates for the extremal eigenvalues of the Schur complements arising in saddle point problems of the form (1).
-If AA is SPD, then the Schur complement $S=BA^{-1}B^{t}$ has the structure of an auxiliary space preconditioner [29, 40], where $A^{-1}$ corresponds to an iterative method defined on the auxiliary space $V$, and $B$ corresponds to the transfer operator mapping the auxiliary space $V$ to the original space $W$.
-A similar argument also applies in the case when AA is semi-SPD.
+If $A$ is SPD, then the Schur complement $S=BA^{-1}B^{t}$ has the structure of an auxiliary space preconditioner [31, 42], where $A^{-1}$ corresponds to an iterative method defined on the auxiliary space $V$, and $B$ corresponds to the transfer operator mapping the auxiliary space $V$ to the original space $W$.
+A similar argument also applies in the case when $A$ is semi-SPD.
 Thus, the Schur complement $S$ can be analyzed by means of the recently proposed auxiliary space theory [35], which provides sharp estimates for iterative methods for both SPD and semi-SPD systems in a unified framework.
-Estimates for the extremal eigenvalues of the Schur complement when AA is SPD are presented in Theorem 4.1, which are sharp in the sense that they are expressed as identities.
+Estimates for the extremal eigenvalues of the Schur complement when $A$ is SPD are presented in Theorem 4.1, which are sharp in the sense that they are expressed as identities.
 Sharp estimates for the Schur complement preconditioned by a general SPD preconditioner are given in Theorem 4.5.
@@
 While our estimates follow directly from elementary linear algebra within the auxiliary space framework, the existing results required more involved arguments.
-Finally, our results extend naturally to the case when AA is semi-SPD, as shown in Theorems 4.9 and 4.10.
+Finally, our results extend naturally to the case when $A$ is semi-SPD, as shown in Theorems 4.9 and 4.10.
 To demonstrate the usefulness of the proposed estimates for the Schur complement, we present several applications from scientific computing.
@@
 Let $V$ and $W$ be finite-dimensional vector spaces equipped with inner products $(\cdot,\cdot)$ and the associated norms $\|\cdot\|$.
-For any subspace $S\subset V$, we denote by $S^{\perp}$ its orthogonal complement.
-For a linear operator $B\colon V\to W$, the null space and the range are denoted by $\mathcal{N}(B)$ and $\mathcal{R}(B)$, respectively, and the adjoint of $B$ is denoted by $B^{t}\colon W\to V$.
+For any subspace $X\subset V$, we denote by $X^{\perp}$ its orthogonal complement.
+For a linear operator $L\colon V\to W$, the null space and the range are denoted by $\mathcal{N}(L)$ and $\mathcal{R}(L)$, respectively, and the adjoint of $L$ is denoted by $L^{t}\colon W\to V$.
 Following standard convention in the literature (see, e.g., [41]), we write $x\lesssim y$, or equivalently $y\gtrsim x$, to mean that there exists a constant $C>0$, independent of the relevant parameters, such that $x\leq Cy$.
@@
 In particular, we present sharp estimates for certain linear operators defined in terms of auxiliary spaces.
-We begin with the auxiliary space lemma [33, Lemma 3.1], which is closely related to the fictitious space method [29] and the auxiliary space method [40].
+We begin with the auxiliary space lemma [35, Lemma 3.1], which is closely related to the fictitious space method [31] and the auxiliary space method [42].
 This lemma plays an important role in the analysis developed in this paper.
@@
 Let $V$ and $\undertilde{V}$ be finite-dimensional vector spaces and let $\Pi\colon\undertilde{V}\to V$ be a surjective linear operator.
-Let $\undertilde{B}\colon\undertilde{V}\to\undertilde{V}$ be a SPD linear operator, and define
-(3) $B=\Pi\undertilde{B}\Pi^{t}\colon V\to V.$
-Then $B$ is SPD. Moreover, it satisfies
-$(B^{-1}v,v)=\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{B}^{-1}\undertilde{v},\undertilde{v}),\quad v\in V.$
-Now, let $A\colon V\to V$ and $B\colon V\to V$ be SPD linear operators, and consider the situation where $B$ is used as a preconditioner for the linear system associated with AA.
-In this case, the convergence rate of iterative methods can be estimated in terms of the condition number of the preconditioned operator BABA, given by
-$\kappa(BA)=\frac{\lambda_{\max}(BA)}{\lambda_{\min}(BA)},$
+Let $\undertilde{M}\colon\undertilde{V}\to\undertilde{V}$ be an SPD linear operator, and define
+(3) $M=\Pi\undertilde{M}\Pi^{t}\colon V\to V.$
+Then $M$ is SPD. Moreover, it satisfies
+$(M^{-1}v,v)=\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{M}^{-1}\undertilde{v},\undertilde{v}),\quad v\in V.$
+Now, let $L\colon V\to V$ and $M\colon V\to V$ be SPD linear operators, and consider the situation where $M$ is used as a preconditioner for the linear system associated with $L$.
+In this case, the convergence rate of iterative methods can be estimated in terms of the condition number of the preconditioned operator $ML$, given by
+$\kappa(ML)=\frac{\lambda_{\max}(ML)}{\lambda_{\min}(ML)},$
 where $\lambda_{\min}$ and $\lambda_{\max}$ denote the minimum and maximum eigenvalues, respectively.
-The following theorem, introduced in [33, Theorem 3.6] and proved using Lemma 2.1, provides sharp estimates for the extremal eigenvalues of BABA when the preconditioner $B$ is defined by (3) for some SPD linear operator $\undertilde{B}\colon\undertilde{V}\to\undertilde{V}$ and a surjective linear operator $\Pi\colon\undertilde{V}\to V$.
+The following theorem, introduced in [35, Theorem 3.6] and proved using Lemma 2.1, provides sharp estimates for the extremal eigenvalues of $ML$ when the preconditioner $M$ is defined by (3) for some SPD linear operator $\undertilde{M}\colon\undertilde{V}\to\undertilde{V}$ and a surjective linear operator $\Pi\colon\undertilde{V}\to V$.
 Here, $\undertilde{V}$ is a finite-dimensional vector space, referred to as an auxiliary space.
-Thus, the preconditioner $B$ is constructed in terms of the operator $\undertilde{B}$ defined on the auxiliary space $\undertilde{V}$.
-Theorem $2.$
-Let $A\colon V\to V$ be a SPD linear operator, and let $B\colon V\to V$ be given in (3). Then we have
-$\lambda_{\min}(BA)$ $=\left(\sup_{v\in V,\ \|v\|_{A}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{B}^{-1}\undertilde{v},\undertilde{v})\right)^{-1},$
-$\lambda_{\max}(BA)$ $=\left(\inf_{v\in V,\ \|v\|_{A}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{B}^{-1}\undertilde{v},\undertilde{v})\right)^{-1}.$
-The result in Theorem 2.2 can be extended to more general situations, for instance when AA is only semi-SPD, $B$ is a general linear operator, and $\undertilde{B}$ is semi-SPD.
+Thus, the preconditioner $M$ is constructed in terms of the operator $\undertilde{M}$ defined on the auxiliary space $\undertilde{V}$.
+Theorem 2.2.
+Let $L\colon V\to V$ be an SPD linear operator, and let $M\colon V\to V$ be given in (3). Then we have
+$\lambda_{\min}(ML)$ $=\left(\sup_{v\in V,\ \|v\|_{L}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{M}^{-1}\undertilde{v},\undertilde{v})\right)^{-1},$
+$\lambda_{\max}(ML)$ $=\left(\inf_{v\in V,\ \|v\|_{L}=1}\inf_{\undertilde{v}\in\undertilde{V},\ \Pi\undertilde{v}=v}(\undertilde{M}^{-1}\undertilde{v},\undertilde{v})\right)^{-1}.$
+The result in Theorem 2.2 can be extended to more general situations, for instance when $L$ is only semi-SPD, $M$ is a general linear operator, and $\undertilde{M}$ is semi-SPD.
 Further details on these extensions can be found in [35].
@@
 3.2 Elimination of the primal variable
-From the saddle point system (1), one can eliminate the primal variable $u$ and obtain a reduced system formulated solely in terms of the dual variable pp.
-Assume first that AA is SPD.
+From the saddle point system (1), one can eliminate the primal variable $u$ and obtain a reduced system formulated solely in terms of the dual variable $p$.
+Assume first that $A$ is SPD.
 From the first equation in (1), we have
@@
 In (7), $S$ is called the Schur complement.
-Because AA is SPD and $B$ is surjective, $S$ is of the form (3).
+Because $A$ is SPD and $B$ is surjective, $S$ is of the form (3).
 In particular, by Lemma 2.1, we deduce that $S$ is SPD.
 Hence, the saddle point system (1) is equivalent to the SPD linear system (7) for $p$, followed by a back substitution to recover $u$.
-The case when AA is only semi-SPD is more delicate, since (6) is not valid.
+The case when $A$ is only semi-SPD is more delicate, since (6) is not valid.
 Nevertheless, elimination of the primal variable $u$ is still possible when full information about the null space $\mathcal{N}(A)$ is available, that is, when an explicit basis of $\mathcal{N}(A)$ is known.
 Such a situation arises, for example, in Neumann boundary value problems for the Poisson equation and linear elasticity [12].
-In the following, we summarize the arguments from [34, Section 2.2] and [37, Section 6.3] concerning the elimination of $u$. Let
+In the following, we summarize the arguments from [36, Section 2.2] and [39, Section 6.3] concerning the elimination of $u$. Let
 $N\colon\mathbb{R}^{\dim\mathcal{N}(A)}\to\mathcal{N}(A)$
 be an injective linear operator with range $\mathcal{N}(A)$.
-Such a map NN can be constructed as the matrix whose columns form a basis of $\mathcal{N}(A)$.
+Such a map $N$ can be constructed as the matrix whose columns form a basis of $\mathcal{N}(A)$.
 From the first equation of (1), compatibility requires $f-B^{t}p\in\mathcal{R}(A)$, which is equivalent to
@@
 (9) $u=A^{+}(f-B^{t}p)+N\xi,$
-for some $\xi\in\mathbb{R}^{\dim\mathcal{N}(A)}$, where $A^{+}\colon V\to V$ denotes a pseudoinverse of AA, i.e., $AA^{+}$ acts as the identity on $\mathcal{R}(A)$.
+for some $\xi\in\mathbb{R}^{\dim\mathcal{N}(A)}$, where $A^{+}\colon V\to V$ denotes a pseudoinverse of $A$, i.e., $AA^{+}$ acts as the identity on $\mathcal{R}(A)$.
 Substituting (9) into the second equation of (1) yields
 (10) $Sp-BN\xi=d,\quad\text{where}\quad S=BA^{+}B^{t}\colon W\to W,\ d=BA^{+}f-g\in W.$
-Note that (10) reduces to (7) when AA is SPD.
+Note that (10) reduces to (7) when $A$ is SPD.
 Now, (8) and (10) form a coupled system in the variables $p$ and $\xi$.
 We next eliminate $\xi$.
-Since $\mathcal{R}(N)=\mathcal{N}(A)$, the condition (5) implies that BNBN is injective.
+Since $\mathcal{R}(N)=\mathcal{N}(A)$, the condition (5) implies that $BN$ is injective.
 Hence the orthogonal projection onto
@@
 (12) $P=I-BN((BN)^{t}BN)^{-1}(BN)^{t}\colon W\to W_{0}.$
-The computation of $P$ requires solving a linear system of $(BN)^{t}BN$, which can be carried out at negligible cost when $\dim\mathcal{N}(A)$ is small.
+The computation of $P$ requires solving a linear system involving $(BN)^{t}BN$, which can be carried out at negligible cost when $\dim\mathcal{N}(A)$ is small.
 We decompose $p$ orthogonally into components in $W_{0}$ and $W_{0}^{\perp}=\mathcal{R}(B|_{\mathcal{N}(A)})$:
@@
 Note that $\eta$ can be computed prior to the main iterations of an iterative method. Finally, substituting (13) into (10) and applying $P$ to both sides yields
-(14) $S_{0}p_{0}=d_{0},\quad\text{where}\quad S_{0}=PSP^{t}\colon W_{0}\to W_{0},\ d_{0}=d-PSBN\eta\in W_{0}.$
-Thus, we arrive at a linear system in the single unknown $p_{0}$, which represents the dual problem in the case when AA is semi-SPD.
+(14) $S_{0}p_{0}=d_{0},\quad\text{where}\quad S_{0}=PSP^{t}\colon W_{0}\to W_{0},\ d_{0}=P(d-SBN\eta)\in W_{0}.$
+Thus, we arrive at a linear system in the single unknown $p_{0}$, which represents the dual problem in the case when $A$ is semi-SPD.
 Proposition 3.1 shows that $S_{0}$ is SPD and independent of the particular choice of the pseudoinverse $A^{+}$ (cf.
-[34, Lemma 2.19]).
+[36, Lemma 2.19]).
 Proposition 3.1.
@@
 This holds because PBPB is surjective and $PB|_{\mathcal{N}(A)}=0$.
-Remark $3.$
+Remark 3.3.
 While we define $P$ as the orthogonal projection onto $W_{0}$ in (12), it is also possible to employ oblique projections, which are often more effective for heterogeneous problems; see [36, 39].
@@
 Here, we discuss iterative methods for solving the saddle point problem (1).
-For simplicity, we assume that AA is SPD.
-Based on the surveys [2, 41], there are three major approaches in iterative methods.
-The first approach is that we apply iterative methods for the SPD dual problems (7) and (14).
+For simplicity, we assume that $A$ is SPD.
+Based on the surveys [2, 43], there are three major approaches to iterative methods.
+The first approach is to apply iterative methods for the SPD dual problem (7) ((14) in the semi-SPD case).
 In this case, we can utilize vast existing results on iterative methods for SPD linear systems.
-See [33] for a systematic approach for design and analysis of iterative methods for SPD linear systems.
+See [35] for a systematic approach to the design and analysis of iterative methods for SPD linear systems.
 The second approach is to apply stationary iterations, such as Uzawa-type methods [5] or the augmented Lagrangian method [13], to (1).
@@
 The third approach is to apply preconditioned Krylov subspace methods, such as MINRES [33], to (1). It was shown in [30] that the block-diagonal preconditioner
-$\mathcal{L}=\begin{bmatrix}A^{-1}&0\\ 0&-S^{-1}\end{bmatrix}$
-is optimal in the sense that the preconditioned operator $\mathcal{L}\mathcal{A}$ has only three distinct eigenvalues, namely 11 and $\tfrac{1}{2}(1\pm\sqrt{5})$.
-Thus, in this case, the central task is the construction of efficient preconditioners for AA and $S$ [2].
+$\mathcal{L}=\begin{bmatrix}A^{-1}&0\\ 0&S^{-1}\end{bmatrix}$
+is optimal in the sense that the preconditioned operator $\mathcal{L}\mathcal{A}$ has only three distinct eigenvalues, namely $1$ and $\tfrac{1}{2}(1\pm\sqrt{5})$.
+Thus, in this case, the central task is the construction of efficient preconditioners for $A$ and $S$ [2].
 In all three approaches, the properties of the Schur complement $S$ are central to both design and analysis of algorithms.
@@
 Although these estimates are derived in a relatively straightforward manner using the auxiliary space theory, they provide bounds that are equivalent to, or even sharper than, existing results, for which more complicated proofs were often required.
-We begin with the case when AA is SPD.
+We begin with the case when $A$ is SPD.
 An obvious but important observation is that the Schur complement $S$ has the same structure as in (3).
-In Theorem 4.1, we establish sharp estimates for the extremal eigenvalues of $S$ defined in (7), obtained as direct consequences of Theorem $2.$
+In Theorem 4.1, we establish sharp estimates for the extremal eigenvalues of $S$ defined in (7), obtained as direct consequences of Theorem 2.2.
 Theorem 4.1.
-In the saddle point problem (1), suppose that AA is SPD. Then the Schur complement $S$ given in (7) satisfies
+In the saddle point problem (1), suppose that $A$ is SPD. Then the Schur complement $S$ given in (7) satisfies
 $\lambda_{\min}(S)=\inf_{0\neq q\in W}\sup_{v\in V,\ Bv=q}\frac{\|q\|^{2}}{(Av,v)},\quad\lambda_{\max}(S)=\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{(Av,v)}.$
@@
 In Theorem 2.2, we set
-$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad A\leftarrow I,\quad B\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{B}\leftarrow A^{-1}.$
+$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad L\leftarrow I,\quad M\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{M}\leftarrow A^{-1}.$
 Then we get
@@
 Moreover, we have
-$\lambda_{\max}(S)=\left(\inf_{0\neq q\in W}\inf_{v\in V,\ Bv=q}\frac{(Av,v)}{\|q\|^{2}}\right)^{-1}=\left(\inf_{0\neq v\in V}\frac{(Av,v)}{\|Bv\|^{2}}\right)^{-1}=\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{(Av,v)},$
+$\lambda_{\max}(S)=\left(\inf_{0\neq q\in W}\inf_{v\in V,\ Bv=q}\frac{(Av,v)}{\|q\|^{2}}\right)^{-1}=\left(\inf_{v\in V,\ Bv\neq 0}\frac{(Av,v)}{\|Bv\|^{2}}\right)^{-1}=\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{(Av,v)},$
 which completes the proof.
-As a corollary of Theorem 4.1, we obtain Corollary 4.3, which appeared in [37, Lemma 9.1] (see also [3]).
+As a corollary of Theorem 4.1, we obtain Corollary 4.3, which appeared in [39, Lemma 9.1] (see also [4]).
 Corollary 4.3.
-In the saddle point problem (1), suppose that AA is SPD. Then the Schur complement $S$ given in (7) satisfies
+In the saddle point problem (1), suppose that $A$ is SPD. Then the Schur complement $S$ given in (7) satisfies
 $\lambda_{\min}(S)\geq\lambda_{\max}(A)^{-1}\|B^{-1}\|^{-2},\quad\lambda_{\max}(S)\leq\lambda_{\min}(A)^{-1}\|B\|^{2}.$
-Proof $4.$
+Proof 4.4.
 By Theorem 4.1, we obtain
 $\lambda_{\min}(S)=\inf_{0\neq q\in W}\sup_{v\in V,\ Bv=q}\frac{\|q\|^{2}}{(Av,v)}\geq\lambda_{\max}(A)^{-1}\inf_{0\neq q\in W}\sup_{v\in V,\ Bv=q}\frac{\|q\|^{2}}{\|v\|^{2}}\\ =\lambda_{\max}(A)^{-1}\left(\sup_{0\neq q\in W}\inf_{v\in V,\ Bv=q}\frac{\|v\|^{2}}{\|q\|^{2}}\right)^{-1}=\lambda_{\max}(A)^{-1}\|B^{-1}\|^{-2},$
-where in the last inequality we use the fact that the minimizer $v\in V$ of $\|v\|^{2}/\|q\|^{2}$ subject to $Bv=q$ must lie in $\mathcal{N}(B)^{\perp}$. Meanwhile, we have
+where in the last equality we use the fact that the minimizer $v\in V$ of $\|v\|^{2}/\|q\|^{2}$ subject to $Bv=q$ must lie in $\mathcal{N}(B)^{\perp}$. Meanwhile, we have
 $\lambda_{\max}(S)=\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{(Av,v)}\leq\lambda_{\min}(A)^{-1}\sup_{0\neq v\in V}\frac{\|Bv\|^{2}}{\|v\|^{2}}=\lambda_{\min}(A)^{-1}\|B\|^{2},$
 which completes the proof.
-Sometimes a preconditioner $L\colon W\to W$ is applied to the Schur complement $S$. Theorem 4.5 provides sharp estimates for the extremal eigenvalues of the preconditioned operator LSLS.
+Sometimes a preconditioner $L\colon W\to W$ is applied to the Schur complement $S$. Theorem 4.5 provides sharp estimates for the extremal eigenvalues of the preconditioned operator $LS$.
 Theorem 4.5.
-In the saddle point problem (1), suppose that AA is SPD. Let $L\colon W\to W$ be a SPD linear operator. Then the Schur complement $S$ given in (7) satisfies
+In the saddle point problem (1), suppose that $A$ is SPD. Let $L\colon W\to W$ be an SPD linear operator. Then the Schur complement $S$ given in (7) satisfies
 $\lambda_{\min}(LS)=\inf_{0\neq q\in W}\sup_{v\in V,\ Bv=q}\frac{(Lq,q)}{(Av,v)},\quad\lambda_{\max}(LS)=\sup_{0\neq v\in V}\frac{(LBv,Bv)}{(Av,v)}.$
@@
 Setting
-$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad A\leftarrow L,\quad B\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{B}\leftarrow A^{-1}$
+$V\leftarrow W,\quad\undertilde{V}\leftarrow V,\quad L\leftarrow L,\quad M\leftarrow S,\quad\Pi\leftarrow B,\quad\undertilde{M}\leftarrow A^{-1}$
 in Theorem 2.2 yields the desired result.
@@
 In this situation, one can construct a special preconditioner $L$ using the right inverse of $B$ as a building block; examples include certain nonoverlapping domain decomposition methods, which we will discuss in Section 7.
-The following corollary, originally stated in [25, Lemma 2] with a rather involved proof, follows directly from Theorem 4.5.
+The following corollary, originally stated in [27, Lemma 2] with a rather involved proof, follows directly from Theorem 4.5.
 Corollary 4.7.
-In the saddle point problem (1), suppose that AA is SPD. Assume further that a linear operator $\bar{B}\colon V\to W$ satisfies $B\bar{B}^{t}=I$, and define a preconditioner $L\colon W\to W$ by
+In the saddle point problem (1), suppose that $A$ is SPD. Assume further that a linear operator $\bar{B}\colon V\to W$ satisfies $B\bar{B}^{t}=I$, and define a preconditioner $L\colon W\to W$ by
 $L=\bar{B}A\bar{B}^{t}.$
@@
 The estimate for the minimum eigenvalue follows by substituting $v=\bar{B}^{t}q$ into the expression in Theorem 4.5, while the estimate for the maximum eigenvalue is an immediate consequence of Theorem 4.5.
-Next, we consider the case when AA is semi-SPD. Recall that the projected Schur complement $S_{0}$ given in (14) satisfies
-$S_{0}=(PB)A^{+}(PB)^{t}.$
+Next, we consider the case when $A$ is semi-SPD. Recall that the projected Schur complement $S_{0}$ given in (14) satisfies
+$S_{0}=PBA^{+}(PB)^{t}.$
 As discussed in the proof of Proposition 3.1, the operator $PB|_{\mathcal{R}(A)}$ is surjective and $\mathcal{R}((PB)^{t})\subset\mathcal{R}(A)$.
-Hence, in view of Proposition 3.1, we may regard $A^{+}$ and PBPB as
-$A^{+}\colon\mathcal{R}(A)\to\mathcal{R}(A),\quad PB\colon\mathcal{R}(A)\to W_{0},$
-where $W_{0}$ was defined in (11). This observation shows that the analysis of the semi-SPD case proceeds in essentially the same way as in the SPD case; see Theorems 4.9 and 4.10.
+Hence, in view of Proposition 3.1, we may regard $A^{+}$ and $PB$ as
+(15) $A^{+}\colon\mathcal{R}(A)\to\mathcal{R}(A),\quad PB\colon\mathcal{R}(A)\to W_{0},$
+where $W_{0}$ was defined in (11).
+This observation shows that the analysis of the semi-SPD case proceeds in essentially the same way as in the SPD case; see Theorems 4.9 and 4.10.
+Consequently, Theorems 4.9 and 4.10 do not depend on a particular choice of the pseudoinverse $A^{+}$ on the whole space $V$, but only on the canonical inverse of $A$ on $\mathcal{R}(A)$.
 Theorem 4.9.
-In the saddle point problem (1), suppose that AA is semi-SPD. Then the projected Schur complement $S_{0}$ given in (14) satisfies
+In the saddle point problem (1), suppose that $A$ is semi-SPD. Then the projected Schur complement $S_{0}$ given in (14) satisfies
 $\lambda_{\min}(S_{0})=\inf_{0\neq q\in W_{0}}\sup_{v\in\mathcal{R}(A),\ PBv=q}\frac{\|q\|^{2}}{(Av,v)},\quad\lambda_{\max}(S_{0})=\sup_{0\neq v\in\mathcal{R}(A)}\frac{\|PBv\|^{2}}{(Av,v)}.$
 Theorem 4.10.
-In the saddle point problem (1), suppose that AA is semi-SPD. Let $L\colon W_{0}\to W_{0}$ be a SPD linear operator. Then the projected Schur complement $S_{0}$ given in (14) satisfies
+In the saddle point problem (1), suppose that $A$ is semi-SPD. Let $L\colon W_{0}\to W_{0}$ be an SPD linear operator. Then the projected Schur complement $S_{0}$ given in (14) satisfies
 $\lambda_{\min}(LS_{0})=\inf_{0\neq q\in W_{0}}\sup_{v\in\mathcal{R}(A),\ PBv=q}\frac{(Lq,q)}{(Av,v)},\quad\lambda_{\max}(LS_{0})=\sup_{0\neq v\in\mathcal{R}(A)}\frac{(LPBv,PBv)}{(Av,v)}.$
@@
 Setting
-$V\leftarrow W_{0},\quad\undertilde{V}\leftarrow\mathcal{R}(A),\quad A\leftarrow L,\quad B\leftarrow S_{0},\quad\Pi\leftarrow PB,\quad\undertilde{B}\leftarrow A^{+}$
+$V\leftarrow W_{0},\quad\undertilde{V}\leftarrow\mathcal{R}(A),\quad L\leftarrow L,\quad M\leftarrow S_{0},\quad\Pi\leftarrow PB,\quad\undertilde{M}\leftarrow A^{+}$
 in Theorem 2.2 yields the desired result.
@@
 Corollary 4.12.
-In the saddle point problem (1), suppose that AA is semi-SPD.
+In the saddle point problem (1), suppose that $A$ is semi-SPD.
 Assume further that a linear operator $\bar{B}\colon V\to W$ satisfies $B\bar{B}^{t}=I$, and define a preconditioner $L\colon W_{0}\to W_{0}$ by
@@
 Then the projected Schur complement $S_{0}$ given in (14) satisfies
-$\lambda_{\min}(LS_{0})\geq 1,\quad\lambda_{\max}(LS_{0})=\sup_{0\neq v\in\mathcal{R}(A)}\frac{(A(P\bar{B})^{t}P\bar{B}v,(P\bar{B})^{t}P\bar{B}v)}{(Av,v)}.$
+$\lambda_{\min}(LS_{0})\geq 1,\quad\lambda_{\max}(LS_{0})=\sup_{0\neq v\in\mathcal{R}(A)}\frac{(A(P\bar{B})^{t}PBv,(P\bar{B})^{t}PBv)}{(Av,v)}.$
 5 Application I: Augmented Lagrangian method
-In the remainder of this paper, we present several examples illustrating how the auxiliary space theory and the estimates for the Schur complements developed in Section 4 can be applied to different applications.
+In the remainder of this paper, we present several examples illustrating how the auxiliary space theory and the estimates for the Schur complements developed in Section 4 can be applied to different problems.
 As a first example, we consider the augmented Lagrangian method [13], which is a stationary iterative method for solving the saddle point problem (1).
 In the augmented Lagrangian method, given the current iterate $(u^{m},p^{m})$, the next iterate $(u^{m+1},p^{m+1})$ is obtained as
-(15) $\begin{aligned} u^{m+1}&=(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon B^{t}g-B^{t}p^{m}),\\ p^{m+1}&=p^{m}-\epsilon^{-1}(g-Bu^{m+1}),\end{aligned}\quad m\geq 0.$
+(16) $\begin{aligned} u^{m+1}&=(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon^{-1}B^{t}g-B^{t}p^{m}),\\ p^{m+1}&=p^{m}-\epsilon^{-1}(g-Bu^{m+1}),\end{aligned}\quad m\geq 0.$
 It is well-known (see, e.g., [2]) that the augmented Lagrangian method (16) is equivalent to the Richardson iteration with step size $\epsilon^{-1}$ applied to the linear system
-(16) $S_{\epsilon}p=d_{\epsilon},\quad\text{where}\quad S_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}B^{t},\,d_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon B^{t}g)-g.$
+(17) $S_{\epsilon}p=d_{\epsilon},\text{ where }S_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}B^{t},\,d_{\epsilon}=B(A+\epsilon^{-1}B^{t}B)^{-1}(f+\epsilon^{-1}B^{t}g)-g.$
 Therefore, the analysis of the augmented Lagrangian method reduces to the analysis of the augmented Schur complement $S_{\epsilon}$.
@@
 $(S_{\epsilon}^{-1}q,q)=\inf_{v\in V,\ Bv=q}((A+\epsilon^{-1}B^{t}B)v,v)\\ =\inf_{v\in V,\ Bv=q}(Av,v)+\epsilon^{-1}(q,q)=(S^{-1}q,q)+\epsilon^{-1}(q,q).$
-Thus we obtain the identity (cf. [18, equation (3.4)]):
-$S_{\epsilon}^{-1}=S^{-1}+\epsilon^{-1}I.$
+Thus we obtain the identity (cf. [20, equation (3.4)]):
+(18) $S_{\epsilon}^{-1}=S^{-1}+\epsilon^{-1}I.$
 Consequently, the extremal eigenvalues of $S_{\epsilon}$ are
@@
 Namely, the convergence rate of (16) becomes arbitrarily fast as $\epsilon$ becomes small.
-This result can be regarded as a refined version of [21, Lemma 2.1], with much simpler proof.
-We note that such arbitrarily fast convergence of the augmented Lagrangian method can be extended to a broader class of convex optimization problems; see [32, Appendix A].
+This result can be regarded as a refined version of [23, Lemma 2.1], with much simpler proof.
+We note that such arbitrarily fast convergence of the augmented Lagrangian method can be extended to a broader class of convex optimization problems; see [34, Appendix A].
 On the other hand, if we apply the conjugate gradient method to solve (17), then the convergence rate depends on the condition number $\kappa(S_{\epsilon})$, which satisfies
@@
 This result both refines and agrees with [19, 20], where the condition number of the augmented Lagrangian method applied to nonoverlapping domain decomposition methods was analyzed.
+Remark 5.1.
+Alternative derivations of the identity (18) can be found in the literature. For example, the well-known Sherman–Morrison–Woodbury formula [15, equation (2.1.4)],
+$(A+UCV)^{-1}=A^{-1}-A^{-1}U(C^{-1}+VA^{-1}U)^{-1}VA^{-1},$
+where AA, $C$, $U$, and $V$ are conformable matrices, can be used to derive (18); see, for example, [3, Lemma 4.1] and [20, equation (3.4)].
 6 Application II: Mixed finite element methods
@@
 (23) $B=-M_{W}\operatorname{div},$
-where $M_{W}\colon W\to W$ is the mass matrix on $W$ defined analogously to $M_{V}$. Moreover, the mass matrices $M_{V}$ and $M_{W}$ satisfy the scaling relations (cf. [37, Section 3.4])
+where $M_{W}\colon W\to W$ is the mass matrix on $W$ defined analogously to $M_{V}$. Moreover, the mass matrices $M_{V}$ and $M_{W}$ satisfy the scaling relations (cf. [39, Section 3.4])
 (24) $(M_{V}v,v)\eqsim h^{d}\|v\|^{2},\quad(M_{W}q,q)\eqsim h^{d}\|q\|^{2},\quad v\in V,\ q\in W,$
@@
 […]
-where the last inequality follows from the standard inverse inequality (see, e.g., [30, Lemma 2.2]). In conclusion, we obtain
+where the last inequality follows from the standard inverse inequality (see, e.g., [32, Lemma 2.2]). In conclusion, we obtain
 $\kappa(S)\lesssim h^{-2}.$
@@
 6.2 Stokes equations
-Let $\Omega$ be a bounded polygonal domain in $\mathbb{R}^{d}$ ($d=2,3$). The incompressible Stokes equations with the homogeneous Dirichlet boundary condition are written as
+Let $\Omega$ be a bounded polyhedral domain in $\mathbb{R}^{d}$ ($d=2,3$). The incompressible Stokes equations with the homogeneous Dirichlet boundary condition are written as
 (26) $-\Delta u+\nabla p$ $=f\quad$ $\text{in }\Omega,$
@@
 A mixed finite element method for solving (26) is obtained by replacing $H_{0}^{1}(\Omega)^{d}$ and $L_{0}^{2}(\Omega)$ with suitable finite element spaces $V$ and $W$.
-Then the corresponding algebraic system becomes (1), where the matrices AA, $B$ and the vector $f$ are defined by
+Then the corresponding algebraic system becomes (1), where the matrices $A$, $B$ and the vector $f$ are defined by
 $(Av,w)=\int_{\Omega}\nabla v\cdot\nabla w\,dx,\ (Bv,q)=-\int_{\Omega}q\operatorname{div}v\,dx,\ (f,v)=\int_{\Omega}f\cdot v\,dx,\quad v,w\in V,\ q\in W.$
@@
 The problem of interest is
-(29) $\min_{v=(v_{j})_{j=1}^{J}\in V}\left\{\frac{1}{2}(Sv,v)-(f,v)\right\}\quad\text{subject to}\quad Bv=0,$
-where the matrix $S$ and the vector $f$ are defined by
-(30) $(Sv,w)=\sum_{j=1}^{J}\int_{\Omega_{j}}\nabla\mathcal{H}_{j}v_{j}\cdot\nabla\mathcal{H}_{j}w_{j}\,dx,\qquad(f,v)=\sum_{j=1}^{J}\int_{\Omega_{j}}f\,\mathcal{H}_{j}v_{j}\,dx,\quad v,w\in V,$
-$\mathcal{H}_{j}$ denotes the discrete harmonic extension in $\Omega_{j}$ associated with $\mathcal{T}_{h}^{j}$ (see [37, 43] for details), $B$ is a full-rank matrix with entries 0 and $\pm 1$ enforcing continuity along $\Gamma$, and $(\cdot,\cdot)$ denotes the Euclidean inner product.
+(31) $\min_{v=(v_{j})_{j=1}^{J}\in V}\left\{\frac{1}{2}(Av,v)-(f,v)\right\}\quad\text{subject to}\quad Bv=0,$
+where the matrix $A$ and the vector $f$ are defined by
+(32) $(Av,w)=\sum_{j=1}^{J}\int_{\Omega_{j}}\nabla\mathcal{H}_{j}v_{j}\cdot\nabla\mathcal{H}_{j}w_{j}\,dx,\qquad(f,v)=\sum_{j=1}^{J}\int_{\Omega_{j}}f\,\mathcal{H}_{j}v_{j}\,dx,\quad v,w\in V,$
+where $\mathcal{H}_{j}$ denotes the discrete harmonic extension in $\Omega_{j}$ associated with $\mathcal{T}_{h}^{j}$ (see [39, 45] for details), $B$ is a full-rank matrix with entries 0 and $\pm 1$ enforcing continuity along $\Gamma$, and $(\cdot,\cdot)$ denotes the Euclidean inner product.
 It is straightforward to observe that (31) is equivalent to the restriction of the finite element discretization of (30) on the global mesh $\mathcal{T}_{h}$ to the interface $\Gamma$.
 By enforcing the constraint $Bv=0$ in (31) through the introduction of a Lagrange multiplier $\lambda\in W$, where $W$ denotes the vector space of Lagrange multipliers, we obtain the following saddle point problem, which is of the form (1):
-(31) $\begin{bmatrix}S&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}f\\ 0\end{bmatrix}$.
-Note that $S$ is semi-SPD, owing to subdomains $\Omega_{j}$ that do not intersect $\partial\Omega$.
+(33) $\begin{bmatrix}A&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}f\\ 0\end{bmatrix}$.
+Note that $A$ is semi-SPD, owing to subdomains $\Omega_{j}$ that do not intersect $\partial\Omega$.
 Finally, by eliminating $u$ in (33) as in Section 4, we obtain the following SPD linear system corresponding to (14):
-(32) $F_{0}\lambda_{0}=d_{0},\quad\text{where}\quad F_{0}=PBS^{+}(PB)^{t}\colon W_{0}\to W_{0},\ d_{0}\in W_{0},$
+(34) $S_{0}\lambda_{0}=d_{0},\quad\text{where}\quad S_{0}=PBA^{+}(PB)^{t}\colon W_{0}\to W_{0},\ d_{0}\in W_{0},$
 and $W_{0}$ and $P$ are defined in the same way as in (11) and (12), respectively.
-Now we estimate the condition number of $F_{0}$ in (32) using the proposed framework.
-We begin by summarizing some useful properties of the matrices $S$ and $B$.
-The matrix $S$ satisfies the following bounds, which combine the well-known estimate for Schur complements in the finite element method (see, e.g., [34, Lemma 1.59]) with standard scaling arguments [37, Section 3.4]:
-(33) $\lambda_{\min}(S)\gtrsim\frac{h}{H},\quad\lambda_{\max}(S)\lesssim 1,$
-where $\lambda_{\min}$ denotes the minimum nonzero eigenvalue.
-Furthermore, as shown in [10, Equation (22)], since the matrix $B$ consists only of entries 0 and $\pm 1$, an elementary calculation gives
+Now we estimate the condition number of $S_{0}$ in (34) using the proposed framework.
+We begin by summarizing some useful properties of the matrices $A$ and $B$.
+We have the following estimates for the extremal eigenvalues of $A$ (cf.
+[11, equations (29) and (30)]).
+Lemma 7.1.
+The matrix AA defined in (32) satisfies
+(35) $\lambda_{\min}(A)\gtrsim\frac{h}{H},\quad\lambda_{\max}(A)\lesssim 1,$
+where $\lambda_{\min}$ denotes the smallest nonzero eigenvalue.
+Proof 7.2.
+For any $v=(v_{j})_{j=1}^{J}\in\mathcal{R}(A)$, we have
+[…]
+[…]
+[…]
+which completes the proof.
+Remark 7.3.
+Let $\hat{V}$ be a subspace of $V$ consisting of functions that are continuous across $\Gamma$.
+Then the restriction $A|_{\hat{V}}$ is SPD, since continuity across $\Gamma$ eliminates the piecewise constant kernel modes associated with floating subdomains.
+For $A|_{\hat{V}}$, we cannot apply the Poincaré inequality on each $\partial\Omega_{j}$ as in Lemma 7.1; instead, we obtain the following weaker estimate for the minimum eigenvalue (see, e.g., [39, Lemma 4.11] and [45, Lemma 3.1]), derived using the global Poincaré inequality on $\Omega$:
+$(Av,v)\gtrsim H\sum_{j=1}^{J}\|v_{j}\|_{L^{2}(\partial\Omega_{j})}^{2}\eqsim Hh(v,v),\quad v=(v_{j})_{j=1}^{J}\in\hat{V}$.
+In addition, as shown in [11, equation (22)], since the matrix $B$ consists only of entries 0 and $\pm 1$, an elementary calculation gives
 (36) $\lambda_{\min}(BB^{t})\gtrsim 1,\quad\lambda_{\max}(BB^{t})=\lambda_{\max}(B^{t}B)\lesssim 1.$
-By Theorem 4.9, the minimum eigenvalue of $F_{0}$ is estimated as
+By Theorem 4.9, the minimum eigenvalue of $S_{0}$ is estimated as
 […]
-where in the second inequality we substituted $v=B^{t}(BB^{t})^{-1}\mu$, which satisfies $PBv=\mu$. On the other hand, the maximum eigenvalue of $F_{0}$ is estimated as
+where, in the second inequality, we substitute
+$v=(PB)^{t}(PB(PB)^{t})^{-1}\mu\in\mathcal{R}(A),$
+which satisfies $PBv=\mu$. Here, the operators $(PB)^{t}$ and $PB(PB)^{t}$ are regarded as
+$(PB)^{t}\colon W_{0}\to\mathcal{R}(A),\quad PB(PB)^{t}\colon W_{0}\to W_{0},$
+in view of (15). On the other hand, the maximum eigenvalue of $S_{0}$ is estimated as
 […]
 Consequently, we get
-$\kappa(F_{0})\lesssim\frac{H}{h}.$
-This result agrees with [10, Theorem 3.2], where the condition number of $F_{0}$ was first estimated.
+$\kappa(S_{0})\lesssim\frac{H}{h}.$
+This result agrees with [11, Theorem 3.2], where the condition number of $S_{0}$ was first estimated.
 Next, we consider the following scaled Dirichlet preconditioner:
-$L_{\mathrm{FETI}}=PB_{D}S(PB_{D})^{t},$
+$L_{\mathrm{FETI}}=PB_{D}A(PB_{D})^{t},$
 where $B_{D}=D_{\mathrm{FETI}}B$ is a scaled jump matrix [38, 39], with $D_{\mathrm{FETI}}$ a block-diagonal scaling matrix chosen such that $BB_{D}^{t}=I$.
@@
 In this case, we can apply Corollary 4.12 to estimate the extremal eigenvalues of the preconditioned operator $L_{\mathrm{FETI}}S_{0}$.
-The only technical ingredient we require is the well-known vertex–edge lemma [43, Section 4.3], stated in Lemma 7.1.
+The only technical ingredient we require is the well-known vertex–edge lemma [45, Section 4.3], stated in Lemma 7.4.
 In what follows, edges are understood as relatively open subsets, i.e., without their endpoints.
@@
     If $x$ is a vertex of $\partial\Omega_{j}$, then we have
-    $\|I_{x}^{0}v\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)\|v\|_{H^{\frac{1}{2}}(\partial\Omega)}^{2},\quad v\in V_{j}.$
+    $\|I_{x}^{0}v_{j}\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)\|v_{j}\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2},\quad v_{j}\in V_{j}.$
   2. (b)
     If $e$ is an edge of $\partial\Omega_{j}$, then we have
-    $\|I_{e}^{0}v\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)^{2}\|v\|_{H^{\frac{1}{2}}(\partial\Omega)}^{2},\quad v\in V_{j}.$
+    $\|I_{e}^{0}v_{j}\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\lesssim\left(1+\log\frac{H}{h}\right)^{2}\|v_{j}\|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2},\quad v_{j}\in V_{j}.$
 Now we are ready to analyze the preconditioned operator $L_{\mathrm{FETI}}S_{0}$. By Corollary 4.12, we have
 $\lambda_{\min}(L_{\mathrm{FETI}}S_{0})\geq 1.$
-To estimate the maximum eigenvalue of $L_{\mathrm{FETI}}S_{0}$, Corollary 4.12 implies that it suffices to bound $(S(P\bar{B})^{t}PBv,(P\bar{B})^{t}PBv)$ for any $v\in\mathcal{R}(S)$.
-Take any $v\in\mathcal{R}(S)$ and choose $\phi\in\mathcal{N}(S)$ such that $v+\phi\in W_{0}$.
+To estimate the maximum eigenvalue of $L_{\mathrm{FETI}}S_{0}$, Corollary 4.12 with $\bar{B}=B_{D}$ implies that it suffices to bound $(A(PB_{D})^{t}PBv,(PB_{D})^{t}PBv)$ for any $v\in\mathcal{R}(A)$.
+Take any $v\in\mathcal{R}(A)$ and choose $\phi\in\mathcal{N}(A)$ such that $B(v+\phi)\in W_{0}$.
 Then we have
@@
 […]
-where the first inequality follows from the definitions of $B$ and $\bar{B}$ together with the triangle inequality, and the second inequality follows from Lemma 7.1.
+where the first inequality follows from the definitions of $B$ and $B_{D}$ together with the triangle inequality, and the second inequality follows from Lemma 7.4.
 Here we require mild geometric assumptions on the domain decomposition $\{\Omega_{j}\}_{j=1}^{J}$, namely that the numbers of vertices and edges of the subdomains are uniformly bounded and that the number of subdomains sharing a common vertex is also uniformly bounded.
 Hence, by Corollary 4.12, we deduce that
-$\lambda_{\max}(L_{\mathrm{FETI}}F_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
-In summary, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{FETI}}F_{0}$ (cf. [37, Theorem 6.15]):
-$\kappa(L_{\mathrm{FETI}}F_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
-We note that the unscaled Dirichlet preconditioner, namely $(PB)S(PB)^{t}$, was first proposed in [10] and analyzed in [26].
+$\lambda_{\max}(L_{\mathrm{FETI}}S_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
+In summary, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{FETI}}S_{0}$ (cf. [39, Theorem 6.15]):
+$\kappa(L_{\mathrm{FETI}}S_{0})\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
+We note that the unscaled Dirichlet preconditioner, namely $PBA(PB)^{t}$, was first proposed in [11] and analyzed in [28].
 In the unscaled case, as discussed in [28], the minimum eigenvalue admits a weaker estimate in certain scenarios.
@@
 Namely, we consider the following formulation:
-(37) $\min_{v=(v_{j})_{j=1}^{J}\in\tilde{V}}\left\{\frac{1}{2}(\tilde{S}v,v)-(f,v)\right\},\quad\text{subject to}\quad Bv=0,$
-where the matrix $\tilde{S}$ and the vector $\tilde{f}$ are defined analogously to (30); see [16, 37] for details.
+(39) $\min_{v=(v_{j})_{j=1}^{J}\in\tilde{V}}\left\{\frac{1}{2}(\tilde{A}v,v)-(\tilde{f},v)\right\},\quad\text{subject to}\quad Bv=0,$
+where the matrix $\tilde{A}$ and the vector $\tilde{f}$ are defined analogously to (32); see [18, 39] for details.
 Here $B$ is a full-rank matrix with entries 0 and $\pm 1$ that enforces continuity along the interior of subdomain edges, and $(\cdot,\cdot)$ denotes the Euclidean inner product.
 By enforcing the constraint $Bv=0$ in (39) through the introduction of a Lagrange multiplier $\lambda\in W$, where $W$ denotes the vector space of Lagrange multipliers, we obtain the saddle point problem
-(38) $\begin{bmatrix}\tilde{S}&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}f\\ 0\end{bmatrix}$.
-Different from the FETI formulation (31), in (38) the matrix $\tilde{S}$ is SPD. Therefore, by eliminating $u$ in (38), we obtain the following Schur complement system corresponding to (7):
-(39) $F\lambda=d,\quad\text{where}\quad F=B\tilde{S}^{-1}B^{t},\quad d=\tilde{S}^{-1}f.$
+(40) $\begin{bmatrix}\tilde{A}&B^{t}\\ B&0\end{bmatrix}\begin{bmatrix}u\\ \lambda\end{bmatrix}=\begin{bmatrix}\tilde{f}\\ 0\end{bmatrix}$.
+In contrast to the FETI formulation (33), in (40) the matrix $\tilde{A}$ is SPD. Therefore, by eliminating $u$ in (40), we obtain the following Schur complement system corresponding to (7):
+(41) $S\lambda=d,\quad\text{where}\quad S=B\tilde{A}^{-1}B^{t},\quad d=B\tilde{A}^{-1}\tilde{f}$.
 In (39), since the matrix $B$ enforces continuity only at nodal points located in the interior of subdomain edges, each nodal point is affected by $B$ only once.
-Hence, we have the following identity [27, Equation (3.2)]:
+Hence, we have the following identity [29, equation (3.2)]:
 (42) $BB^{t}=2I.$
@@
 As FETI-DP involves functions that are continuous at subdomain corners, we require certain Poincaré-type inequalities associated with subdomain corners in order to analyze FETI-DP.
-In Lemma 7.2, which summarizes results from [19, Lemma 2] and [27, Lemma 5.1], $L^{2}$ and $H^{\frac{1}{2}}$ estimates are provided for functions vanishing at subdomain corners.
+In Lemma 7.5, which summarizes results from [21, Lemma 2] and [29, Lemma 5.1], $L^{2}$ and $H^{\frac{1}{2}}$ estimates are provided for functions vanishing at subdomain corners.
 In what follows, let $I_{H}$ denote the linear coarse interpolation operator associated with the coarse triangulation $\mathcal{T}_{H}$.
@@
 where $I_{e}^{0}$ was defined in (37).
-Now, we estimate the extremal eigenvalues of $F$ given in (39). By Theorem 4.1, we obtain
+Now, we estimate the extremal eigenvalues of $S$ given in (41). By Theorem 4.1, we obtain
 […]
@@
 where the last inequality is due to Lemma 7.5. Hence, Theorem 4.1 implies that
-$\lambda_{\max}(F)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right).$
+$\lambda_{\max}(S)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right).$
 Consequently, we deduce that
-$\kappa(F)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right),$
-which agrees with the existing result in [19, Proposition 3].
+$\kappa(S)\lesssim\frac{H}{h}\left(1+\log\frac{H}{h}\right),$
+which agrees with the existing result in [21, Proposition 3].
 Next, we consider the following Dirichlet preconditioner:
-$L_{\mathrm{DP}}=\bar{B}\tilde{S}\bar{B}^{t},$
-By Corollary 4.7, we have $\lambda_{\min}(L_{\mathrm{DP}}F)\geq 1$. Moreover, for any $v=(v_{j})_{j=1}^{J}\in\tilde{V}$, it follows that
-$\begin{split}(\tilde{S}\bar{B}^{t}Bv,\bar{B}^{t}Bv)&=(\tilde{S}\bar{B}^{t}B(v-I_{H}v),\bar{B}^{t}B(v-I_{H}v))\\ &=\sum_{j=1}^{J}|\bar{B}^{t}B(v-I_{H}v)|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\sum_{j=1}^{J}\sum_{e\text{:~edge of }\Omega_{j}}\left|I_{e}^{0}(v_{j}-I_{H}v_{j})\right|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\left(1+\log\frac{H}{h}\right)^{2}\sum_{j=1}^{J}|v_{j}|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &=\left(1+\log\frac{H}{h}\right)^{2}(\tilde{S}v,v),\end{split}$
+$L_{\mathrm{DP}}=\bar{B}\tilde{A}\bar{B}^{t},$
+By Corollary 4.7, we have $\lambda_{\min}(L_{\mathrm{DP}}S)\geq 1$. Moreover, for any $v=(v_{j})_{j=1}^{J}\in\tilde{V}$, it follows that
+$\begin{split}(\tilde{A}\bar{B}^{t}Bv,\bar{B}^{t}Bv)&=(\tilde{A}\bar{B}^{t}B(v-I_{H}v),\bar{B}^{t}B(v-I_{H}v))\\ &\eqsim\sum_{j=1}^{J}|\bar{B}^{t}B(v-I_{H}v)|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\sum_{j=1}^{J}\sum_{e\text{:~edge of }\Omega_{j}}\left|I_{e}^{0}(v_{j}-I_{H}v_{j})\right|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\lesssim\left(1+\log\frac{H}{h}\right)^{2}\sum_{j=1}^{J}|v_{j}|_{H^{\frac{1}{2}}(\partial\Omega_{j})}^{2}\\ &\eqsim\left(1+\log\frac{H}{h}\right)^{2}(\tilde{A}v,v),\end{split}$
 where the first inequality follows from the definitions of $B$ and $\bar{B}$ together with the triangle inequality, and the second inequality is due to Lemma 7.5. This implies
-$\lambda_{\max}(L_{\mathrm{DP}}F)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
-Finally, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{DP}}F$ (cf. [27, Theorem 5.2]):
-$\kappa(L_{\mathrm{DP}}F)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
+$\lambda_{\max}(L_{\mathrm{DP}}S)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
+Finally, we obtain the following condition number bound for the preconditioned operator $L_{\mathrm{DP}}S$ (cf. [29, Theorem 5.2]):
+$\kappa(L_{\mathrm{DP}}S)\lesssim\left(1+\log\frac{H}{h}\right)^{2}.$
 The above proof can be extended to the cases of three dimensions or heterogeneous coefficients [18, 39] without major difficulties.
@@
 Two closely related nonoverlapping domain decomposition methods to FETI and FETI-DP are BDD (balancing domain decomposition) [14, 25] and BDDC (balancing domain decomposition by constraints) [9, 14].
-It has been shown in the literature [6, 22, 24, 25, 36] that FETI/FETI-DP with scaled Dirichlet preconditioners, and BDD/BDDC are spectrally equivalent in the sense that their spectra agree except for the eigenvalues 0 and 11.
+It has been shown in the literature [7, 24, 26, 27, 38] that FETI/FETI-DP with scaled Dirichlet preconditioners, and BDD/BDDC are spectrally equivalent in the sense that their spectra agree except for the eigenvalues 0 and $1$.
 In fact, BDD and BDDC can be analyzed within our framework (Theorems 4.5 and 4.10) by essentially the same arguments used for FETI and FETI-DP.
@@
 In this paper, we have presented sharp estimates for the Schur complements arising in saddle point problems, by utilizing the auxiliary space theory introduced in [35].
-We also provided various examples, such as the augmented Lagrangian method, mixed finite element methods, and nonoverlapping domain decomposition methods, which highlight the usefulness of our results.
+We have also provided various examples, such as the augmented Lagrangian method, mixed finite element methods, and nonoverlapping domain decomposition methods, which highlight the usefulness of our results.
 In particular, the proposed framework simplifies the analysis compared to existing approaches by minimizing the algebraic effort required.
@@
 The second remark concerns the extension of the proposed framework to a broader range of problems.
-As mentioned at several points in this paper (see, e.g., Remark 3.4), some parts of the contents of this paper can already be extended to more general classes of convex optimization problems, although a full generalization is highly nontrivial.
+As mentioned at several points in this paper (see, e.g., Remark 3.4), some of the results in this paper can already be extended to more general classes of convex optimization problems, although a full generalization is highly nontrivial.
 We note that there has recently been significant research on extending the theory of iterative methods for linear systems to convex optimization problems; see [16] and the references therein.